### v1

# MathGAP: Out-of-Distribution Evaluation on Problems with Arbitrarily Complex Proofs

Abstract
Large language models (LLMs) can solve arithmetic word problems with high accuracy, but little is known about how well they generalize to problems that are more complex than the ones on which they have been trained. Empirical investigations of such questions are impeded by two major flaws of current evaluations: (i) much of the evaluation data is contaminated, in the sense that it has already been seen during training, and (ii) benchmark datasets do not capture how problem proofs may be arbitrarily complex in various ways. As a step towards addressing these issues, we present a framework for evaluating LLMs on problems that have arbitrarily complex arithmetic proofs, called MathGAP. MathGAP generates problems that follow fixed proof specifications—along with chain-of-thought reasoning annotations—enabling systematic studies on generalization with respect to arithmetic proof complexity. We apply MathGAP to analyze how in-context learning interacts with generalization to problems that have more complex proofs. We find that among the models tested, most show a significant decrease in performance as proofs get deeper and wider. This effect is more pronounced in complex, nonlinear proof structures, which are challenging even for GPT-4o. Surprisingly, providing in-context examples from the same distribution as the test set is not always beneficial for performance. In particular, zero-shot prompting as well as demonstrating a diverse range of examples that are less complex than the test data sometimes yield similar or higher accuracies.
1 Introduction
High performance on reasoning benchmarks are often taken as evidence that transformer-based large language models (LLMs) are able to “reason”. However, many current evaluations are unreliable since it is likely that the problems they contain are present in the model’s training data (Sainz et al., 2023; Deng et al., 2024; Zhang et al., 2024). Moreover, most evaluations fail to capture that reasoning problems can be arbitrarily complex, through composition of subproofs and use of multiple rules of inference. High accuracy on a specific set of problems in a benchmark dataset is therefore not sufficient to conclude that LLMs can generalize to more complex, unseen problems. To obtain a more appropriate empirical measure of reasoning ability, one must evaluate on data that is not present in any benchmark dataset, containing proofs that are more complex than those used in training and context window alike.
Math word problems (MWPs) are one of the most frequently used testbeds for LLM reasoning. Yet there exist few, if any, systematic analyses on out-of-distribution (OOD) generalization in regards to proof complexity that avoid contamination in this domain. We suspect that this is in part due to a lack of formalism for characterizing proof complexity, which we argue is necessary for making precise claims about generalization behavior for reasoning problems.
Figure 1: We propose MathGAP, an evaluation framework for arithmetic reasoning in which LLMs are tested on problems with proofs of arbitrary complexity. This diagram shows how problems and CoT solution annotations are generated under our formalism. The complete list of logical forms and inference rules that we consider in our experiments are given in Tables 1 and 3.1, respectively.
In this paper, we present a framework for evaluating Mathematical Generalization on Arithmetic Proofs—MathGAP. First, we discuss how to characterize solutions to MWPs in terms of proof trees, in which the nodes are labeled with logical statements about the problem under a world-model framework (Opedal et al., 2023) and the edges are induced by proof steps on such logical statements. Doing so reveals several ways of characterizing a problem’s complexity based on its proof tree, e.g., (non)linearity, depth, width, and the ordering of nodes (all defined in § 3.2). With this formalism, we develop a method to generate problems with specified proof tree characteristics. Because this method is based on proof trees, it also enables rich annotations in the form of ground-truth reasoning traces. More specifically, traversing the proof trees in post order, we obtain chain-of-thought (CoT; Wei et al., 2022) reasoning traces in an automated manner. Fig. 1 illustrates the generation method, showing a nonlinearauthor=bernhard,color=white!40,size=,fancyline,caption=,]does everybody know what nonlinear means in this context, or should this be defined? Could mention that this will be defined below, at the end of Sec 3  andreas: good point, added above. I am now worried that (non)linearity is not a standard term here. Maybe we should rebrand it to something like “proof tree topology”? Would be more general as well proof tree with depth $2$ and width $5$, in which the ordering of the sentences is given by traversing the leaf nodes in left-to-right order. This generation method forms the backbone of MathGAP, our framework for studying OOD generalization gaps on arithmetic proofs. MathGAP avoids data contamination since the distribution of generated test problems can always be different from those in training or in context—by design. When the performance on problems at one level of complexity hits saturation, we can flexibly generate a new set of problems which are even more complex. This is in spirit similar to Dynabench (Kiela et al., 2021) but, importantly, does not require human annotators.
MathGAP opens up the possibility for several new empirical investigations. [^1] Code to generate problems with MathGAP will be publicly released with the final version. In the present study, we perform a systematic analysis on whether LLMs can use simple examples in context to solve more complex ones at inference. To this end, we consider two different distributions of in-context CoT demonstrations: examples with only the simplest possible proof trees, and a range of examples with varying complexity, akin to curriculum learning (Bengio et al., 2009). We then evaluate on test sets containing problems that are increasingly more complex than the in-context examples. The performance is compared to two baselines of in-context demonstrations, a prompt that has in-distribution examples with respect to the test set, and a zero-shot prompt. Our study includes Mixtral-8x7B (Jiang et al., 2024), Llama3-8B, Llama3-70B (Llama Team, 2024), GPT-3.5-Turbo, GPT-4o (OpenAI, 2024), and a few additional evaluations on o1-preview. We present several findings:
  • As expected, there is consistent decrease in performance as proof depth and width increase. The decrease is particularly notable for nonlinear problems, for which we observe a general inability of the evaluated LLMs to solve complex problems, even for GPT-4o.
  • Generalization across orderings appears to be different, however. We conduct experiments where we move sentences to the front of the problem, and find that the problem is easiest for LLMs if the sentence is moved from near the beginning or end, rather than from the middle of the problem. This is surprising since one would expect a monotonically decreasing relationship between accuracy and the distance of the movement.
  • We see no clear relationship between in-context distribution and performance. Contrary to our expectations, in-distribution examples are not always beneficial for performance, as compared to zero shot. Demonstrating a range of examples with varying complexityauthor=bernhard,color=white!40,size=,fancyline,caption=,]sounds like in-context curriculum learning?  andreas: yes true, I’ve added some references to that. The difference is that we randomly sample the ordering of the examples; this difference is explained in § 5 is usually preferable to demonstrating simple examples, suggesting that LLMs benefit from a diverse prompt.
  • The o1-preview model outperforms the other models on nonlinear problems. Yet, we are able to create a test set of which it solves only 5%—demonstrating that MathGAP is future-proof.
2 Related Work
Data contamination.
Machine learning, broadly, is the study of algorithms that can learn from data, and generalize to unseen data. Machine learning models must therefore be evaluated on unseen test sets. As it turns out, modern-day LLMs are not being properly evaluated in this regard: Much of the data they are evaluated on has already been used during model training, both labeled (Dodge et al., 2021; Deng et al., 2024) and unlabeled (Elazar et al., 2023). Researchers have raised serious concerns about these issues (Jacovi et al., 2023; Sainz et al., 2023). For arithmetic reasoning specifically, Zhang et al. (2024) recently found evidence that the widely used GSM8k benchmark (Cobbe et al., 2021) is contaminated, and presented evaluations on a new dataset which is not publicly released. MathGAP mitigates data contamination by creating new, more complex synthetic test sets. author=mrinmaya,color=pink!40,size=,fancyline,caption=,]I am less sure about this. Maybe we should just say that we can use MATHGap to always create new test problems  andreas: Yeah I guess that’s a stronger argument, edited. However I think we should refrain from publishing our test data though. Instead only publish the data-generating code. Do you disagree?  abu: oh interesting, so keep the random seed for our test data hidden? i suppose people can still reproduce our results to an extent, with sampling noise  andreas: yes exactly, just so that models don’t overfit to our particular test examples
Generalization on reasoning tasks.
Generalization to data from a different distribution (i.e., OOD data) is a core problem in machine learning, and has been studied extensively also in the context of LLM reasoning (Schwarzschild et al., 2021; Razeghi et al., 2022; An et al., 2023; Kudo et al., 2023; Liu et al., 2023; Saparov & He, 2023; Zhang et al., 2023; Thomm et al., 2024). While LLMs’ performance on OOD data can be improved by various techniques (Anil et al., 2022; Hu et al., 2024), generalization to problems of arbitrary complexity under a set of known inference rules is still an open problem. For instance, Dziri et al. (2023) demonstrate that the accuracy of transformer-based LLMs on compositional generalization rapidly approaches zero as the complexity of the problem increases—regardless of in-context guidance. Saparov et al. (2023) focus on theorem proving for logical reasoning, presenting a systematic analysis on generalization ability of LLMs with respect to proof size and diversity of inference rules. Our paper contributes to this literature by providing an evaluation framework to systematically study OOD generalization in regards to arithmetic proof complexity.
Evaluation on MWP benchmarks.
It is common to treat MWPs as a testbed for studying reasoning abilities of LLMs (Cobbe et al., 2021; Patel et al., 2021; Fu et al., 2023; Shakarian et al., 2023; Stolfo et al., 2023; Zong & Krishnamachari, 2023); such problems are useful because solving them requires several distinct skills, yet, they remain conceptually simple (Stern, 1993). However, we are not aware of many studies that investigate generalization in regards to problem complexity in this domain. Our study presented in § 5 relates to Hase et al. (2024), who fine-tune LLMs on easy problems and evaluate them on harder ones at inference. Their complexity metrics focus on problem length, but do not capture how a longer reasoning problem gets more complex (i.e., the shape of the proof tree). Moreover, they evaluate on GSM8k, which is contaminated. Mishra et al. (2024) proposed a method that uses GPT-4 to generate new synthetic problems from symbolic specifications. Their approach is scalable, but using an LLM to generate problems may (i) introduce bias (Boguraev et al., 2024; Opedal et al., 2024) and (ii) produce problems that are unfaithful to their specifications. Mirzadeh et al. (2024) recently performed evaluations on a version of GSM8k in which the variable names and numbers have been altered, and found that LLMs are surprisingly sensitive to such substitutions. MathGAP is more general; while it allows for such substitutions, our main goal is to vary the proof structures.
Table 1: Examples of logical forms. A logical form consists of a predicate and a set of properties which is specific to that predicate, given here in separate columns. Each logical form induces a set of natural language sentences which preserve its semantics.
Logical Form Example Templates Example Sentences
Predicate (abbr.) Properties
container (agent=a, [a] has [q] [u]s of [k] [e]s. Alice has 5 kilograms of red apples.
(cont) quantity=q,
entity=e, [a] owns [q] [u]s of [k] [e]s. Alice owns 5 kilograms of red apples.
attribute=k,
unit=u)
comparison (agentA=a, [b] has [q] fewer [e]s than [a]. Bob has 3 fewer apples than Alice.
(comp) agentB=b,
quantity=q, [a] has [q] more [e]s than [b]. Alice has 3 more apples than Bob.
entity=e)
transfer (receiver_agent=b, [a] gave [b] [q] [e]s. Alice gave Bob 3 apples.
sender_agent=a,
quantity=q, [b] got [q] more [e]s from [a]. Bob got 3 more apples from Alice.
entity=e)
partwhole (whole_agent=$\hskip 1.00006pt\land\hskip 1.00006pt_{i=1}^{n}\texttt{a}_{i}$, [$\texttt{a}_{1}$], …, and [$\texttt{a}_{n}$] combine the [f]s that they have. Alice and Bob combine the fruits that they have.
part_agent1=a1,
part_agentn=an,
whole_entity=f,
part_entity=e)
comp-eq (agentA=a, The number of [e]s that [c] has more than [d] is equal to the difference between the number of [e]s that [a] and [b] have. The number of apples that Charlie has more than David is equal to the difference between the number of apples that Alice and Bob have.
agentB=b,
agentC=c,
agentD=d,
entity=e)
3 A Formal Treatment of Math Word Problems
We aim to generate new problems of arbitrary complexity, so we require a formalism for characterizing complexity in a precise manner. First, we explain how the semantics expressed in math word problems can be written as sequences of logical forms (§ 3.1). We then show how to apply inference rules on these logical forms to deduce new ones (§ 3.2). This leads to a view of problem solutions as proof trees, the structure of which can be used to characterize the reasoning required to solve the problem.
3.1 Math word problems as logical forms
Each math word problem is represented as a sequence of logical forms under the formalism from Opedal et al. (2023; 2024). A logical form is a statement about a truth in the world that is being described by the problem, representing some arithmetic relationship between the number of entities possessed by one or several agents. Such statements can refer to the number of entities an agent has, how many more entities an agent has relative to another, the action of an agent giving their entities to another agent, etc. Formally, a logical form represents the semantics of a single sentence, and consists of a predicate that takes a set of properties as arguments. Every logical form has an agent property, an entity property and a quantity property. Other, optional, properties include unit and attribute. We define a world model as a sequence of logical forms that align with the sentences of a problem text, thus, representing the semantics of the entire problem. See Fig. 5 for the world model representation of the problem from Fig. 1.
For a given predicate and set of properties, we write predicate(property1=x, property2=y, ...) to express a logical form. Property names are, however, often omitted for brevity, as in predicate(x, y, ...). There are two types of predicates: (i) those that represent ownership over some entity by an agent, called container (abbreviated cont), and (ii) those that represent arithmetic relationships between two sets of properties. The relationship predicates correspond to arithmetic concepts and are based on a taxonomy from the learning sciences (Riley et al., 1983); we use transfer , comparison (abbreviated comp), partwhole and comp-eq. [^2] The transfer and part-whole concepts are usually called change and combine, respectively, in learning science literature (Nesher et al., 1982). The comp-eq predicate was introduced in this paper in order to construct nonlinear problems of arbitrary length. Table 1 gives concrete example sentences for logical forms with each of these predicates. In words, a cont logical form expresses the fact that an agent owns a particular entity in some quantity (e.g., “Alice has 5 apples”). A comp logical form expresses how much more some agent owns of a particular entity than some other agent (e.g., “Bob has 3 fewer apples than Alice”). A transfer logical form expresses the transfer of a certain quantity of an entity from one agent to another (e.g., “Alice gave Bob 3 apples”). A partwhole logical form expresses a superset relation between the combined entities of several agents (e.g., Alice and Bob combine their fruits). Finally, a comp-eq logical form expresses that the quantity of a particular comp between two agents is equal to that of a comp between two other agents (e.g., “The difference between the number of apples that Alice and Bob have is the same as the difference between the number of apples that Charlie and David have”).
The logical forms in the world model of a math problem are separated into the body and question. The question is a single logical form which appears at the end of the world model. It is characterized by having a placeholder variable instead of an explicit quantity, which corresponds to the correct answer of the problem. For example, the interrogative sentence “How many apples does Alice have?” is represented by the question cont(Alice, q, apples), where q is a placeholder variable. The body consists of all logical forms in the world model which are not questions. The answer to a problem is the logical form declared by the question with the correct numerical quantity substituted into its placeholder variable, although we will sometimes refer to only the correct numerical quantity as the answer as well.
Table 2: Inference rule templates (left) with corresponding examples of rule applications in natural language (right). The variables a, b, c, d refer to agents, q, q1, q2 refer to quantities, and e, f refer to entities. Attribute and unit properties are excluded here, but analogous rules exist in which they are present. All inference rules are commutative except the transfer rule; see footnote 3. Rule applications of an inference rule with premises $L_{1},L_{2},\dots,L_{N}$ and consequent $L$ are written as $L^{\prime}_{1}.L^{\prime}_{2}.\,\dots\,.L^{\prime}_{N}.\vdash L^{\prime}.$, where $L^{\prime}$ is a natural language expression for the logical form $L$. Axioms are formed from the facts stated in the problem, as illustrated in Fig. 1.
Inference Rules Example Sentences
[…]
3.2 Chain-of-thought solutions as proof trees
Solving math word problems is a form of deductive reasoning, where the correct answer follows necessarily from what is stated in the text through a combination of rules of arithmetic and world knowledge. We use such derivations, called proofs, to characterize the structure of problems. The proofs in our setup are analogous to those found in other deductive systems, e.g., of first-order logic.author=andreas,color=blue!40,size=,fancyline,caption=,]This could be a good place to cite other NLP papers which have formalized reasoning through deduction systems. Any suggestions? abu:there is a lot we could cite here; maybe going all the way back to Montague, surveys of formal semantics, categorial grammar (where parsing itself is formalized as a reasoning problem)
[…]
A math word problem can be solved by applying inference rules until the answer of the problem has been proved. Formally, a proof tree is a rooted ordered tree where each node is labeled with a logical form $L$, which is the conclusion of a proof step with premises $L_{1},\dots,L_{N}$, matching the labels of the child nodes in the same order. Leaf nodes are thus labeled with axioms. A proof of a particular problem is a proof tree where the labels on the leaf nodes are logical forms in the body of the problem, and the label at the root is the answer to the problem. The tree shown in Fig. 1 illustrates a proof of the problem given in the same figure.
Proof complexity.
We seek a systematic way to analyze whether LLMs can generalize from simple to more complex proofs. The definition of proof given above gives rise to several ways of characterizing reasoning, out of which we consider four: linearity, depth, width, and ordering. These are defined below.
We say that a proof tree is linear if every one of its inference rules takes at most one premise that is not an axiom, and nonlinear otherwise. In other words, every step in the solution to a linear problem only uses at most one fact not directly present in the problem. The proof tree given in Fig. 1 is nonlinear because the proof step that deduces the answer uses two premises that are not themselves axioms.author=andreas,color=blue!40,size=,fancyline,caption=,]We should give an example of a linear problem as well so that this is easier to understand abu: agreed, maybe in appendix? The depth of a proof tree is defined as its height, i.e., the maximum path length from the root node to any leaf node. Next, the width of a proof tree is the number of leaf nodes, or axioms, that it contains. The proof tree in Fig. 1 thus has depth 2 and width 5. In words, this means that the answer is two proof steps “away” from any of the axioms, and that the problem has five (declarative) sentences. [^4] In this paper, the width of the proof tree matches the number of declarative sentences (i.e., all sentences except the question) in the corresponding problem since we restrict the logical forms and sentences to be one-to-one. In general, that need not be the case since a single sentence could be written as a conjunction of several logical forms. One could also imagine problems that have irrelevant sentences, which do not represent any logical form at all. Lastly, we define the canonical ordering to be the ordering of leaf nodes as visited left-to-right. Note that the sentences in the natural language annotation from Fig. 1 follows the canonical ordering of the proof tree. Informally, this is the easy way of ordering the sentences in the problem, since it matches the ordering of the proof steps.author=andreas,color=blue!40,size=,fancyline,caption=,]Would be great if we could find support for this statement in cogsci literature However, there exist multiple alternative orderings in general. In particular, given a proof tree that only contains commutative inference rules (see footnote 3), all orderings of the leaf nodes are valid. For instance, swapping the ordering of the first two leaf nodes in the tree in Fig. 1 yields a problem with the same proof.
4 Evaluation Framework
Equipped with the formalism explained in § 3, we propose an evaluation framework which tests LLMs on MWPs of arbitrary proof complexity, called MathGAP. In § 4.1, we explain the generation method behind MathGAP. We then outline how this method can be used for OOD evaluation in § 4.2.
4.1 Generation Method
We propose a method for synthetic generation of problems and their CoT solution explanations. It consists of three high-level steps: (i) sample a proof tree, (ii) map the logical forms at the leaf nodes into a natural language problem, and (iii) map the proof steps into a CoT solution. [^5] Jin et al. (2023) presents a related method for generating causal reasoning problems. Fig. 1 gives an illustration of these steps; note that the agent properties are abbreviated for brevity.
In detail, the procedure is as follows: We first sample a logical form at the root of the proof tree. We then sample a specific inference rule as well as its premises. We recursively repeat this procedure for each premise until a predetermined stopping criterion has been reached, resulting in a proof tree. The stopping criterion is user-specified; it could be, e.g., when the tree has reached a certain depth or width. The categorical properties for the logical forms (i.e., agent, entity, attribute, and unit) are sampled uniformly at random from handwritten vocabularies. The quantities are sampled uniformly at random from a user-specified range; in our study presented in § 5 we use $2{-}20$.
We then traverse the leaves of the proof tree in some order and convert each into a sentence via a natural language template, forming the text corresponding to the body of the problem. The default order follows the proof tree’s canonical ordering, but other orders are possible as well. The sentences are sampled uniformly at random from a predefined set of templates; some example templates are given in Table 1. The question is converted from the logical form at the root of the proof tree, and is sampled from a set of interrogative sentences.
In the final step, we generate the CoT annotation for the problem. We first obtain a sequential ordering of the CoT proof steps by visiting the nodes in the proof tree with a post-order traversal. [^6] Note that the choice of traversing the nodes in post order corresponds to applying each inference rule as soon as all of its premises are available. In addition, post-order traversal visits the axioms in the canonical ordering. Alternative orders of tree traversal would violate at least one of these constraints.author=andreas,color=blue!40,size=,fancyline,caption=,]If we want to comment on similarity to the ordering of human reasoning etc we should look into literature on that. Start with the figural effect reference by Eisape et al. (2024) and the proof ordering references by Boguraev et al. (2024) We then convert each proof step into a sentence via a natural language template. The templates used for the axioms are identical to those used in the body of the problem text. In other words, the proof steps corresponding to axioms simply repeat the relevant part of the problem text. For the other proof steps, we have templates which are designed to explain the computation step that is performed. Fig. 1 illustrates the mapping between nodes in the proof tree and sentences in the CoT annotation. For instance, the third sentence in Fig. 1 explains the conclusion cont(Lucy, 27, apple), which is deduced from the two leftmost nodes in the tree: “So Lucy has $17+10=27$ apples”.
Our pipeline is similar to the one proposed by Opedal et al. (2024), except that we use a recursive procedure to generate each proof step starting at the root, instead of a sequential one. This is a more general approach which also enables generation of nonlinear proofs. [^7] The comp-eq rule enables the generation of nonlinear proof trees with arbitrary depth and is not included in the list of relations from Opedal et al. (2023).author=andreas,color=blue!40,size=,fancyline,caption=,]We need to explain the nuance of building nonlinear proofs in an appendix. We can create a nonlinear ones of depth two with only comp , but for deeper than that we need comp-eq. In the experiments we sample depth 2 examples of both comp and comp-eq.
4.2 MathGAP: Mathematical Generalization on Arithmetic Proofs
With the generation method just described, we can generate new synthetic problems where we control for the structure of the proof and its complexity. MathGAP can be used for various forms of systematic studies on arithmetic reasoning; here, we focus on OOD generalization. Specifically, we can select a family of problems of interest (e.g., linear problems under comp inference rules) and generate a train-test split of problems for which the problems in the test set are more complex than those in the training set (e.g., containing problems with deeper proofs).
The generation method can be flexibly adapted to fit specific studies. One can add new logical forms and inference rules, restrict the choice of existing ones during sampling, or create logical forms for irrelevant sentences (Shi et al., 2023). One can also vary the text templates for the logical forms and inference rules, the vocabularies for agents, entities, attributes and units, and the range of numbers from which the quantities are sampled. In addition, one may increase linguistic diversityauthor=andreas,color=blue!40,size=,fancyline,caption=,]There must be a better, more linguistically correct term for this by incorporating a last step in which a language model paraphrases the templated texts, or, alternatively, use a procedure like the one proposed by Mishra et al. (2024). In our study we do not perform such paraphrasing, since we want to guarantee that the sentences are faithful to the proof trees from which they are generated. However, future studies using MathGAP may choose to prioritize differently.
5 Experiments
author=andreas,color=blue!40,size=,fancyline,caption=,]Remind the reader somewhere here of why our evaluation framework is neat. We could have done an analysis on, say, GSM8k, but we prefer our synthetic data so that we can (i) avoid data contamination and (ii) control the proof tree characteristics
We apply MathGAP to study how LLMs generalize from proof demonstrations of simple problems in context to solve problems of higher complexity at inference. We test OOD generalization through several sets of experiments: depth generalization for linear problems (§ 5.1), width generalization for linear problems (§ 5.2), depth generalization for nonlinear problems (§ 5.3), and generalization to permutations (§ 5.4). We are also interested in how the choice of the distribution of problems provided in context affects performance, which is often not obvious (Min et al., 2022; Turpin et al., 2023).
Experimental setup.
For each set of experiments, the general experimental setup is as follows: We generate five test sets of different degrees of complexity with 400 problems in each. We then generate model predictions for the problems in these test sets under four different prompts: (i) A zero-shot baseline prompt, which contains only the test problem; (ii) A prompt of primitive examples, in which all problems contain only one proof step of the same inference rule that is used in the test problem; (iii) A prompt of examples within a range of complexities, all of which are simpler than the test problem;author=abu,color=orange!40,size=,fancyline,caption=,]maybe we should format these settings differently so that it’s clearer that we will refer to them later in this section; either by bolding or maybe by naming them “Prompt (i):” (iv) An in-distribution baseline prompt, containing examples that are of the same complexity as the test problem.
[…]
Responses are generated using greedy decoding and a maximum length of 4,096 tokens. Model predictions are obtained by extracting the last number occurring in the model’s generated output. We report answer accuracies and compute 95% confidence intervals using bootstrap sampling (Efron, 1992). We evaluate the following models on all test sets: Mixtral-8x7B (Jiang et al., 2024), Llama3 with 8B and 70B parameters (Llama Team, 2024), GPT-3.5 Turbo and GPT-4o (OpenAI, 2024). [^10] The maximum context windows for these models are 65k, 8k, 8k, 16k, 128k, respectively. Combined with a maximum number of output tokens of 4,096, the models have a theoretically sufficient context size to solve all test problems. The largest test problems (nonlinear depth 6) have around 2.5k tokens, including CoT annotation. For the OpenAI models we used “gpt-3.5-turbo-0125” and “gpt-4o-2024-05-13” as the model ids for all experiments.
Figure 2: Answer accuracies for generalization to increasing depth and width for linear problems, across models and in-context distributions. Depth is increased using inference rules involving comp and transfer . Width is increased using partwhole.
5.1 Linear Depth Generalization
Problem sets.
We consider linear problems using both comp and transfer .author=andreas,color=blue!40,size=,fancyline,caption=,]We should give example proof trees for all our settings in appendix The five test sets consist of problems with depths between $6{-}10$. As such, the examples provided in the range of setting (iii) described above have depths between $1{-}5$. We give an example of a problem with depth $6$ in § B.1.
Results.
We illustrate the results in the top and middle rows of Fig. 2. As expected, we observe a decreasing trend in performance as depth increases, for all in-context distributions. Curiously, GPT-3.5-Turbo is the only model for which adding in-context examples consistently leads to improvements. For the majority of cases, performance tends to be larger when providing a range of examples of varying complexity as compared to only primitive examples. The ranged OOD cases often show comparable performance to that of in-distribution context, suggesting that these models benefit from demonstrations on how problems of varying complexity are solved. Llama3-8B is an outlier in that its solving accuracy for in-distribution context is (substantially) lower as compared to other prompts. When inspecting the model outputs, we see that the model often fails to answer the inference question, recognizing instead that it has been provided a sequence of in-context examples. This could potentially be a consequence of how Llama3 is trained, using smaller context windows for the initial stages of pretraining (Llama Team, 2024). Finally, we note that GPT-4o and the larger Llama3 model show close to perfect performance across all test sets.
5.2 Linear Width Generalization
Problem sets.
To test generalization to greater proof width, we use partwhole problems with a fixed depth of $1$. The five test sets have widths between $6{-}10$ and the width of the problems in setting (iii) described above have widths between $1{-}5$.
Results.
The results are shown in the bottom row of Fig. 2. For GPT-3.5 and Mixtral-8x7B, we observe that providing in-context examples improves accuracy over zero-shot in most cases. Like for linear depth generalization, it seems that having a range of OOD examples of varying complexity is superior to only including primitive examples. Llama3-8B exhibits worse performance with in-distribution context here as well. In general for the Llama models, adding OOD in-context examples usually leads to lower performance over zero-shot. Apart from GPT-3.5-Turbo, the performance is generally lower as compared to a deep problem with the same number of sentences (§ 5.1), often with a sharper decreasing trend. Such is the case even for Llama3 70B, which performed almost perfectly in the linear depth setting. This would suggest that it is more difficult to generalize in terms of width than in terms of depth. Again, GPT-4o demonstrates near perfect performance across all test sets.
Figure 3: Answer accuracies for generalization to increasing depth and width for nonlinear problems, across models and in-context distributions. Depth is increased using inference rules involving comp and comp-eq . The in-distribution contexts were too large to fit in the prompts for the Llama3 models and GPT-3.5 Turbo at some of the higher depths. App. C shows additional results on o1-preview.
5.3 Nonlinear Depth Generalization
Problem sets.
We generate test datasets containing nonlinear problems with comp and comp-eq , having the same form as the problem in Fig. 1. Another example is shown in § B.2. We generate a test set for each depth between $3{-}6$. Thus, the ranged OOD context setting contains examples of depths between $1{-}2$.author=andreas,color=blue!40,size=,fancyline,caption=,]Mention somewhere how depth interacts with width for these problem types
Results.
The results are shown in Fig. 3. All models exhibit a performance trend that tends to zero as depth increases, albeit at different rates. For Mixtral-8x7B, Llama3-8B, and GPT-3.5 Turbo, the performance decreases rapidly across all context modes. These models benefit from in-distribution contexts for low depths, and in some cases, from a range of OOD examples. Llama3-70B and GPT-4o are more robust to this type of OOD distribution shift, [^11] Note that GPT-4o performs better on depth 6 problems than depth 5 problems under the in-distribution prompt. Although we provided the same model id as in the other cases—which according to the documentation points to the original “gpt-4o-2024-05-13” snapshot—this particular test was performed at a later point in time. but for the deepest problems, their performance tends to zero as well. Moreover, it is worth noting that, for many cases, providing in-context examples does not significantly improve performance over zero-shot. On the other hand, for lower depths it is usually beneficial to provide in-distribution contexts. We perform an additional analysis on o1-preview in App. C.
5.4 Order Generalization
Problem sets.
LLMs are known to be sensitive to changes in the order of axioms in logical and mathematical reasoning problems (Chen et al., 2024; Eisape et al., 2024). Here, we present a systematic analysis of their sensitivity to a particular form of order permutation. We consider linear, left-leaning proof trees using comp that have depth 5, and generate problems that deviate from the canonical, left-to-right order of the leaf nodes. In particular, we pick one of the leaf nodes to visit first, and then visit the remaining leaf nodes in left-to-right order. This has the effect of moving one of the sentences in the canonically-ordered problem text to the beginning of the problem. We characterize complexity as the distance of the movement, with the hypothesis that a greater movement from a canonical ordering constitutes a more difficult problem.author=andreas,color=blue!40,size=,fancyline,caption=,]Add a figure to the appendix where we show a linear proof tree and number the canonical orderings and give examples of movements we consider. Similar style as Fig. 1 but less stuff in it. We create five test sets with movement distances between $1{-}5$. We do not include a primitive OOD in-context distribution, the notion of a primitive problem is ill-defined in this case. The range prompt includes movements of all distances between $1{-}5$ except the one present in the test set.
Figure 4: Answer accuracies for generalization to permutations, across models and in-context distributions. We measure complexity as the distance of the movement of a sentences moved to the beginning of the problem, as compared to its position in the canonical ordering of the problem.
Results.
The results are shown in Fig. 4. First, note that all models show lower performance on perturbed problems as compared to non-perturbed ones (compare to Fig. 2). Second, there is a nonlinear relationship between accuracy and distance of the movement, in contrast to the monotonically decreasing relation we expected. More specifically, the performance is the highest for the problems where the sentence is moved from near the beginning or the end of the problem. [^12] This is consistent with findings on retrieval-augmented generation, for which models are better at using information that occurs near the beginning or end of the prompt (Liu et al., 2024a). In addition, including in-context examples gives a larger boost in performance for short and long movements, as compared to medium-length ones.
6 Conclusion
In this paper, we introduced MathGAP, a framework for evaluating LLMs on math word problems of arbitrary arithmetic proof complexity. MathGAP can flexibly generate synthetic arithmetic word problems with controllable proof tree characteristics. Importantly, this enables studies of OOD generalization that avoid issues of data contamination, since a train-test split where the distributions are different can be created programmatically. In particular, the test set can be made arbitrarily more complex than the training set. We apply MathGAP to study whether LLMs can learn from simple examples in context to solve more complex problems in a test set. All LLMs tested show a decrease in performance as proof complexity increases through depth and/or width. We also find that LLMs are sensitive to the order in which sentences are presented in a particular manner: A problem is harder if a sentence is moved from the middle of the problem, as compared to the beginning or end. We find no clear relationship between the distribution of in-context examples and performance; zero-shot or OOD contexts sometimes yield performance higher than or comparable to in-distribution contexts. Finally, we demonstrate that we can use MathGAP to construct problems that are challenging even for state-of-the-art reasoning models like OpenAI o1—see App. C.
Limitations
Although we make sure to sample from a wide range of agent names, entities, attributes and units, and include multiple templates for the predicates, future work may investigate how our findings generalize to distributions of problems with higher linguistic diversity. It is also beyond the scope of the present study to consider problem texts in languages other than English; however, it is straightforward to extend MathGAP to other languages. In addition, we cannot guarantee that the distributions of our generated problems are different from those of any potential internal data-generating mechanisms from Meta or OpenAI. However, we deem it to be unlikely that they use a similar formalism as we do here, and it gets increasingly unlikely that problems are similar as complexity increases.
While the more complex nonlinear problems are indeed challenging even for the most capable models, we note that the main purpose of our study was not to create a challenging evaluation set for state-of-the-art LLMs. If that is the goal, one could create harder problems by combining several inference rules (e.g., including the rate concept; Opedal et al., 2023), designing a wider variety of proof tree topologies, and presenting the axioms in arbitrary orderings. It is also possible to create more syntactically complex templates and use numbers on which it has been shown to be more challenging for models to perform arithmetic (Razeghi et al., 2022).
Finally, it is hard to make a direct comparison between the difficulty of linear and nonlinear problems, since they use different inference rules. In particular, nonlinear proofs contain comp-eq predicates, which are not present in linear proofs.
Acknowledgments
We thank Alessandro Stolfo, Shehzaad Dhuliawala, Yanick Zengaffinen, and Yilmazcan Ozyurt for useful feedback and discussion. Andreas Opedal acknowledges funding from the Max Planck ETH Center for Learning Systems.
Appendix A Example of a World Model
Fig. 5 gives an example problem text along with its world model.
Figure 5: World model (right) of a math word problem (left). Each sentence in the problem text is represented by a logical form which consists of a predicate with property arguments. The logical forms in the body are used as axioms in the proof of the problem, as shown in Fig. 1. The sentences and logical forms labeled (1) through (5) are the body of the problem, (6) is the question and (7) is the answer.
Appendix B Example Problems
author=yilmazcan,color=white!40,size=,fancyline,caption=,]For each subsection, maybe a small comment about why this problem is depth 6 or 3, or width 6 etc?
In this section we provide example problems for each set of experiments (§§ 5.1, 5.2, 5.3 and 5.4).
B.1 Example linear comparison problem (depth 6)
Jackson has 16 red bottles of soap. Jackson has 10 more red bottles of soap than Abigail. Joseph has 18 more red bottles of soap than Abigail. Joseph has 14 fewer red bottles of soap than James. Michael has 2 more red bottles of soap than James. Ryan has 16 fewer red bottles of soap than Michael. Mia has 10 more red bottles of soap than Ryan. What is the number of red bottles of soap that Mia has?
B.2 Example nonlinear comparison problem (depth 3)
Ella has 11 yellow plates. Ella has 19 fewer yellow plates than Jacob. Evelyn has 16 yellow plates. Daniel has 10 yellow plates. The number of yellow plates that Emma has more than Jacob is the same as the difference between the number of yellow plates that Evelyn has compared to Daniel. Lucy has 2 yellow plates. Amelia has 6 more yellow plates than Lucy. Layla has 17 yellow plates. Layla has 13 fewer yellow plates than Sophia. The number of yellow plates that Emma has more than Henry is the same as the difference between the number of yellow plates that Amelia has compared to Sophia. What is the number of yellow plates that Henry has?
B.3 Example partwhole width problem (width 7)
Emily has 5 apples. Lily has 8 bananas. Abigail has 9 bananas. Benjamin has 11 grapes. Christopher has 20 apples. Mila has 16 grapes. Sophia has 11 watermelons. If everyone sums up the fruits that they have, how many fruits does everybody have in total?
B.4 Permuted problems
Canonically ordered problem text
Nicholas has 19 computers. Lucy has 6 fewer computers than Nicholas. Harper has 6 fewer computers than Lucy. John has 10 more computers than Harper. Abigail has 15 fewer computers than John. Abigail has 12 fewer computers than Logan. How many computers does Logan have in their collection?
Permuted problem text (distance 1)
Lucy has 6 fewer computers than Nicholas. Nicholas has 19 computers. [Original Position] Harper has 6 fewer computers than Lucy. John has 10 more computers than Harper. Abigail has 15 fewer computers than John. Abigail has 12 fewer computers than Logan. How many computers does Logan have in their collection?
Permuted problem text (distance 3)
John has 10 more computers than Harper. Nicholas has 19 computers. Lucy has 6 fewer computers than Nicholas. Harper has 6 fewer computers than Lucy. [Original Position] Abigail has 15 fewer computers than John. Abigail has 12 fewer computers than Logan. How many computers does Logan have in their collection?
Permuted problem text (distance 5)
Abigail has 12 fewer computers than Logan. Nicholas has 19 computers. Lucy has 6 fewer computers than Nicholas. Harper has 6 fewer computers than Lucy. John has 10 more computers than Harper. Abigail has 15 fewer computers than John. [Original Position] How many computers does Logan have in their collection?
Appendix C Additional Analysis on OpenAI o1
Fig. 6 shows additional results on the OpenAI o1-preview model, “o1-preview-2024-09-12”, evaluated on the nonlinear test sets. We only used zero-shot prompting due to the high costs of performing inference with o1-preview. Using a simple, straightforward prompt is also consistent with OpenAI’s current recommendations. We see that the o1 model shows superior performance to all models from the main text. However, as with the other models, the performance decreases as complexity increases.
Importantly, we note that performance can be improved by increasing the limit on the number of output tokens. The results presented in Fig. 6 used a token limit of 4,096, like the experiments in the main paper. We compared the effect of the token limit on an additional test set of 400 nonlinear problems with depth 7. With a token limit of 4,096, the model could only answer $0.25\%$ of the problems correctly. However, with a token limit of 10,000, it answered $76.5\%$ of the problems correctly. These results are noteworthy, and suggests that the new inference technique used by the o1 model is highly effective in decomposing complex problems into smaller steps.
To test the model’s limits, we performed a final evaluation on a subset of 100 of the depth 7 problems where the sentences in each problem were randomly ordered. Recall from § 3.2 that permuting the sentences in random order is valid since all the inference rules used in these problems are commutative. In this experiment, we allowed 25,000 output tokens, which is the current recommendation by OpenAI. On this set, the answer accuracy was only $5\%$. Thus, we conclude that MathGAP is future-proof in the sense that it can generate problems on which current state-of-the-art models fail. We stress, however, that the aim of this paper is not to construct a challenging test set for state-of-the-art models; there are several other ways that MathGAP can be used to create problems that might be even more difficult.
Figure 6: Answer accuracies for generalization to increasing depth and width for nonlinear problems for OpenAI o1-preview under zero-shot prompting.

### v3

@@
 Abstract
-Large language models (LLMs) can solve arithmetic word problems with high accuracy, but little is known about how well they generalize to problems that are more complex than the ones on which they have been trained.
-Empirical investigations of such questions are impeded by two major flaws of current evaluations: (i) much of the evaluation data is contaminated, in the sense that it has already been seen during training, and (ii) benchmark datasets do not capture how problem proofs may be arbitrarily complex in various ways.
-As a step towards addressing these issues, we present a framework for evaluating LLMs on problems that have arbitrarily complex arithmetic proofs, called MathGAP.
-MathGAP generates problems that follow fixed proof specifications—along with chain-of-thought reasoning annotations—enabling systematic studies on generalization with respect to arithmetic proof complexity.
-We apply MathGAP to analyze how in-context learning interacts with generalization to problems that have more complex proofs.
-We find that among the models tested, most show a significant decrease in performance as proofs get deeper and wider.
-This effect is more pronounced in complex, nonlinear proof structures, which are challenging even for GPT-4o.
-Surprisingly, providing in-context examples from the same distribution as the test set is not always beneficial for performance.
-In particular, zero-shot prompting as well as demonstrating a diverse range of examples that are less complex than the test data sometimes yield similar or higher accuracies.
+Large language models (LLMs) can solve arithmetic word problems with high accuracy, but little is known about how well they generalize to more complex problems.
+This is difficult to study, as (i) much of the available evaluation data has already been seen by the most capable models during training, and (ii) existing benchmarks do not capture how problem proofs may be arbitrarily complex in various ways.
+In this paper, we present a data-generation framework for evaluating LLMs on problems with arbitrarily complex arithmetic proofs, called MathGAP.
+MathGAP generates problem statements and chain-of-thought reasoning traces according to specifications about their arithmetic proof structure, enabling systematic studies on easy-to-hard generalization with respect to complexity of proof trees.
+Using MathGAP, we find that LLMs show a significant decrease in performance as proofs get deeper and wider.
+This effect is more pronounced in complex, nonlinear proof structures, which are challenging even for the most capable models.
+The models are also sensitive to simple changes in sentence ordering.
+However, they remain capable of solving some complex problems, suggesting that reasoning generalization is noisy.
+[Uncaptioned image]   https://github.com/eth-lre/mathgap-experiments
 1 Introduction
-High performance on reasoning benchmarks are often taken as evidence that transformer-based large language models (LLMs) are able to “reason”.
-However, many current evaluations are unreliable since it is likely that the problems they contain are present in the model’s training data (Sainz et al., 2023; Deng et al., 2024; Zhang et al., 2024).
-Moreover, most evaluations fail to capture that reasoning problems can be arbitrarily complex, through composition of subproofs and use of multiple rules of inference.
+High performance on reasoning benchmarks is often taken as evidence that transformer-based large language models (LLMs) can “reason”.
+However, many current benchmarks are unreliable, as it is likely that the problems they contain are present in the model training data (Sainz et al., 2023; Deng et al., 2024; Zhang et al., 2024).
+Moreover, most evaluations fail to capture that reasoning problems can be arbitrarily complex through composition of subproofs and use of multiple rules of inference.
 High accuracy on a specific set of problems in a benchmark dataset is therefore not sufficient to conclude that LLMs can generalize to more complex, unseen problems.
-To obtain a more appropriate empirical measure of reasoning ability, one must evaluate on data that is not present in any benchmark dataset, containing proofs that are more complex than those used in training and context window alike.
-Math word problems (MWPs) are one of the most frequently used testbeds for LLM reasoning.
-Yet there exist few, if any, systematic analyses on out-of-distribution (OOD) generalization in regards to proof complexity that avoid contamination in this domain.
-We suspect that this is in part due to a lack of formalism for characterizing proof complexity, which we argue is necessary for making precise claims about generalization behavior for reasoning problems.
+To obtain a more appropriate empirical measure of reasoning ability, one must evaluate on data that is not present in any benchmark dataset, containing proofs that are more complex than the ones the model has already seen.
 Figure 1: We propose MathGAP, an evaluation framework for arithmetic reasoning in which LLMs are tested on problems with proofs of arbitrary complexity.
@@
 The complete list of logical forms and inference rules that we consider in our experiments are given in Tables 1 and 3.1, respectively.
-In this paper, we present a framework for evaluating Mathematical Generalization on Arithmetic Proofs—MathGAP.
-First, we discuss how to characterize solutions to MWPs in terms of proof trees, in which the nodes are labeled with logical statements about the problem under a world-model framework (Opedal et al., 2023) and the edges are induced by proof steps on such logical statements.
-Doing so reveals several ways of characterizing a problem’s complexity based on its proof tree, e.g., (non)linearity, depth, width, and the ordering of nodes (all defined in § 3.2).
+In this paper, we aim to satisfy these desiderata.
+We consider math word problems (MWPs), which are one of the most frequently used testbeds for LLM reasoning, and propose a framework for evaluating Mathematical Generalization on Arithmetic Proofs—MathGAP.
+We start by discussing how solutions to MWPs can be characterized in terms of proof trees, in which nodes are labeled with logical statements about the problem under a world-model framework (Opedal et al., 2023), and the edges are induced by proof steps on such logical statements.
+Doing so reveals several ways of characterizing a problem’s complexity based on its proof tree, e.g., depth, width, shape, and the ordering of nodes (all defined in § 3.2).
 With this formalism, we develop a method to generate problems with specified proof tree characteristics.
@@
 Fig.
-1 illustrates the generation method, showing a nonlinearauthor=bernhard,color=white!40,size=,fancyline,caption=,]does everybody know what nonlinear means in this context, or should this be defined?
-Could mention that this will be defined below, at the end of Sec 3  andreas: good point, added above.
-I am now worried that (non)linearity is not a standard term here.
-Maybe we should rebrand it to something like “proof tree topology”?
-Would be more general as well proof tree with depth $2$ and width $5$, in which the ordering of the sentences is given by traversing the leaf nodes in left-to-right order.
-This generation method forms the backbone of MathGAP, our framework for studying OOD generalization gaps on arithmetic proofs.
-MathGAP avoids data contamination since the distribution of generated test problems can always be different from those in training or in context—by design.
-When the performance on problems at one level of complexity hits saturation, we can flexibly generate a new set of problems which are even more complex.
+1 illustrates the generation method, showing a proof tree having depth $2$, width $5$, what we call nonlinear shape, and where the ordering of the sentences is given by traversing the leaf nodes in left-to-right order.
+This generation method forms the backbone of MathGAP, our framework for studying out-of-distribution (OOD) generalization gaps on arithmetic proofs.
+MathGAP avoids data contamination since the distribution of generated test problems can always be different from those in training or in context, by design.
+When the performance on problems at one level of complexity hits saturation, we can flexibly generate a new set of problems that are even more complex.
 This is in spirit similar to Dynabench (Kiela et al., 2021) but, importantly, does not require human annotators.
-MathGAP opens up the possibility for several new empirical investigations.
-[^1] Code to generate problems with MathGAP will be publicly released with the final version.
-In the present study, we perform a systematic analysis on whether LLMs can use simple examples in context to solve more complex ones at inference.
+MathGAP enables many new empirical investigations.
+In this paper, we perform a systematic analysis on whether LLMs can use simple examples in context to solve more complex ones at inference.
 To this end, we consider two different distributions of in-context CoT demonstrations: examples with only the simplest possible proof trees, and a range of examples with varying complexity, akin to curriculum learning (Bengio et al., 2009).
 We then evaluate on test sets containing problems that are increasingly more complex than the in-context examples.
-The performance is compared to two baselines of in-context demonstrations, a prompt that has in-distribution examples with respect to the test set, and a zero-shot prompt.
-Our study includes Mixtral-8x7B (Jiang et al., 2024), Llama3-8B, Llama3-70B (Llama Team, 2024), GPT-3.5-Turbo, GPT-4o (OpenAI, 2024), and a few additional evaluations on o1-preview.
+The performance is further compared to in-distribution and zero-shot baselines.
+Our study includes Mixtral-8x7B (Jiang et al., 2024a), Llama3-8B, Llama3-70B (Llama Team, 2024), GPT-3.5-Turbo, GPT-4o (OpenAI, 2024), and a few additional evaluations on o1-preview and DeepSeek-R1 (DeepSeek-AI, 2025).
 We present several findings:
-  • As expected, there is consistent decrease in performance as proof depth and width increase.
-The decrease is particularly notable for nonlinear problems, for which we observe a general inability of the evaluated LLMs to solve complex problems, even for GPT-4o.
-  • Generalization across orderings appears to be different, however.
-We conduct experiments where we move sentences to the front of the problem, and find that the problem is easiest for LLMs if the sentence is moved from near the beginning or end, rather than from the middle of the problem.
-This is surprising since one would expect a monotonically decreasing relationship between accuracy and the distance of the movement.
-  • We see no clear relationship between in-context distribution and performance.
-Contrary to our expectations, in-distribution examples are not always beneficial for performance, as compared to zero shot.
-Demonstrating a range of examples with varying complexityauthor=bernhard,color=white!40,size=,fancyline,caption=,]sounds like in-context curriculum learning?
-andreas: yes true, I’ve added some references to that.
-The difference is that we randomly sample the ordering of the examples; this difference is explained in § 5 is usually preferable to demonstrating simple examples, suggesting that LLMs benefit from a diverse prompt.
-  • The o1-preview model outperforms the other models on nonlinear problems. Yet, we are able to create a test set of which it solves only 5%—demonstrating that MathGAP is future-proof.
+  1. i)
+    Generalization to proof width is harder than to proof depth, but there is a steady decrease in performance as proofs get both deeper and wider.
+The performance decrease is particularly notable for the so-called nonlinear problems; solving them requires keeping intermediate steps longer in memory than for linear problems, for which intermediate steps can be consumed immediately (under post-order traversal).
+Even GPT-4o fails to solve the most complex nonlinear problems.
+  2. ii)
+    However, the performance decrease is moderate in most cases. This suggests that the models are able to generalize to some extent, albeit with a noisy form of reasoning (Prabhakar et al., 2024).
+  3. iii)
+    All models are sensitive to a simple change in sentence ordering, in which one sentence is moved to the front of the problem.
+Moreover, we find that the problem is easiest for LLMs if the sentence is moved from the beginning or the end rather than from the middle of the problem.
+This is surprising since one would expect a monotonically decreasing relationship between accuracy and the distance of movement.
+  4. iv)
+    We do not see a consistent advantage to using in-context examples.
+Contrary to our expectations, in-distribution examples are not always beneficial for performance, as compared to zero-shot.
+This suggests that some problems are hard to learn even when the LLM is given demonstrations of similar problems, independently of its generalization ability.
+Demonstrating a range of examples with varying complexity is usually preferable to demonstrating simple examples, suggesting that LLMs benefit from a diverse prompt.
+  5. v)
+    o1 and R1 outperform the others on nonlinear problems. However, we are able to create a test set on which they solve only 5% and 11% respectively—indicating that MathGAP is future-proof.
 2 Related Work
 Data contamination.
-Machine learning, broadly, is the study of algorithms that can learn from data, and generalize to unseen data.
+Machine learning, broadly, is the study of algorithms that can learn from data and generalize to unseen data.
 Machine learning models must therefore be evaluated on unseen test sets.
-As it turns out, modern-day LLMs are not being properly evaluated in this regard: Much of the data they are evaluated on has already been used during model training, both labeled (Dodge et al., 2021; Deng et al., 2024) and unlabeled (Elazar et al., 2023).
-Researchers have raised serious concerns about these issues (Jacovi et al., 2023; Sainz et al., 2023).
+As it turns out, modern-day LLMs are not being properly evaluated in this regard.
+Much of the data they are evaluated on has already been used during model training, both labeled (Dodge et al., 2021; Deng et al., 2024) and unlabeled (Elazar et al., 2024).
+Researchers have raised serious concerns about such data contamination issues (Jacovi et al., 2023; Sainz et al., 2023).
 For arithmetic reasoning specifically, Zhang et al.
-(2024) recently found evidence that the widely used GSM8k benchmark (Cobbe et al., 2021) is contaminated, and presented evaluations on a new dataset which is not publicly released.
+(2024) recently found evidence that the widely used GSM8k benchmark (Cobbe et al., 2021) is contaminated and presented evaluations on a new dataset that is not publicly released.
 MathGAP mitigates data contamination by creating new, more complex synthetic test sets.
-author=mrinmaya,color=pink!40,size=,fancyline,caption=,]I am less sure about this.
-Maybe we should just say that we can use MATHGap to always create new test problems  andreas: Yeah I guess that’s a stronger argument, edited.
-However I think we should refrain from publishing our test data though.
-Instead only publish the data-generating code.
-Do you disagree?
-abu: oh interesting, so keep the random seed for our test data hidden?
-i suppose people can still reproduce our results to an extent, with sampling noise  andreas: yes exactly, just so that models don’t overfit to our particular test examples
 Generalization on reasoning tasks.
-Generalization to data from a different distribution (i.e., OOD data) is a core problem in machine learning, and has been studied extensively also in the context of LLM reasoning (Schwarzschild et al., 2021; Razeghi et al., 2022; An et al., 2023; Kudo et al., 2023; Liu et al., 2023; Saparov & He, 2023; Zhang et al., 2023; Thomm et al., 2024).
-While LLMs’ performance on OOD data can be improved by various techniques (Anil et al., 2022; Hu et al., 2024), generalization to problems of arbitrary complexity under a set of known inference rules is still an open problem.
-For instance, Dziri et al.
-(2023) demonstrate that the accuracy of transformer-based LLMs on compositional generalization rapidly approaches zero as the complexity of the problem increases—regardless of in-context guidance.
-Saparov et al.
-(2023) focus on theorem proving for logical reasoning, presenting a systematic analysis on generalization ability of LLMs with respect to proof size and diversity of inference rules.
-Our paper contributes to this literature by providing an evaluation framework to systematically study OOD generalization in regards to arithmetic proof complexity.
+Generalization to data from a different distribution (i.e., OOD data) is a core problem in machine learning and has also been studied extensively in the context of LLM reasoning (Schwarzschild et al., 2021; An et al., 2023; Kudo et al., 2023; Liu et al., 2023; Saparov & He, 2023; Zhang et al., 2023; Mészáros et al., 2024; Thomm et al., 2024).
+While LLMs’ performance on OOD data can be improved by various techniques (Anil et al., 2022; Borazjanizadeh & Piantadosi, 2024; Hu et al., 2024), generalization to problems of arbitrary complexity under a set of known inference rules is still an open problem.
+For instance, transformer-based LLMs struggle to generalize to longer problem proofs when solving logical reasoning tasks (Dziri et al., 2023; Saparov et al., 2023).
+We contribute to this literature by providing an evaluation framework for OOD generalization to complex arithmetic proofs.
+Our work relates to scalable oversight (Bowman et al., 2022) and easy-to-hard generalization (Burns et al., 2023; Hase et al., 2024; Sun et al., 2024), which can be studied in a principled manner with MathGAP by characterizing what is easy and what is hard.
 Evaluation on MWP benchmarks.
-It is common to treat MWPs as a testbed for studying reasoning abilities of LLMs (Cobbe et al., 2021; Patel et al., 2021; Fu et al., 2023; Shakarian et al., 2023; Stolfo et al., 2023; Zong & Krishnamachari, 2023); such problems are useful because solving them requires several distinct skills, yet, they remain conceptually simple (Stern, 1993).
-However, we are not aware of many studies that investigate generalization in regards to problem complexity in this domain.
-Our study presented in § 5 relates to Hase et al.
-(2024), who fine-tune LLMs on easy problems and evaluate them on harder ones at inference.
-Their complexity metrics focus on problem length, but do not capture how a longer reasoning problem gets more complex (i.e., the shape of the proof tree).
-Moreover, they evaluate on GSM8k, which is contaminated.
-Mishra et al.
-(2024) proposed a method that uses GPT-4 to generate new synthetic problems from symbolic specifications.
-Their approach is scalable, but using an LLM to generate problems may (i) introduce bias (Boguraev et al., 2024; Opedal et al., 2024) and (ii) produce problems that are unfaithful to their specifications.
+MWPs are often treated as a testbed for studying reasoning abilities of LLMs (Patel et al., 2021; Fu et al., 2023; Shakarian et al., 2023; Stolfo et al., 2023; Zong & Krishnamachari, 2023; Huang et al., 2024; Ye et al., 2024; Yu et al., 2024); such problems are interesting because they require several distinct skills to solve, while remaining conceptually simple (Stern, 1993).
+However, we are not aware of much work on generalization in regard to problem complexity in this domain.
+Our study in § 5 is related to Hase et al.
+(2024), who fine-tune LLMs on easy problems and evaluate them on harder ones at inference, using GSM8k and other possibly contaminated datasets.
+Their complexity metrics focus on problem length, but do not capture how a reasoning problem gets more complex (i.e., through the characteristics of the proof tree).
 Mirzadeh et al.
-(2024) recently performed evaluations on a version of GSM8k in which the variable names and numbers have been altered, and found that LLMs are surprisingly sensitive to such substitutions.
-MathGAP is more general; while it allows for such substitutions, our main goal is to vary the proof structures.
-Table 1: Examples of logical forms.
-A logical form consists of a predicate and a set of properties which is specific to that predicate, given here in separate columns.
-Each logical form induces a set of natural language sentences which preserve its semantics.
+(2025) recently performed evaluations on a version of GSM8k in which variable names and numbers have been altered, and found that LLMs are surprisingly sensitive to such substitutions.
+MathGAP is more general; while it allows for such substitutions as well, our main goal is to vary the proof structures.
+3 A Formal Treatment of Math Word Problems
+We aim to generate new problems of arbitrary complexity, so we require a formalism to characterize complexity in a precise manner.
+First, we explain how the semantics expressed in MWPs can be written as sequences of logical forms (§ 3.1).
+We then show how to apply the inference rules on these logical forms to deduce new ones (§ 3.2).
+This leads to a view of problem solutions as proof trees, the structure of which can be used to characterize the reasoning required to solve the problem.
+Table 1: Example logical forms.
+A logical form comprises a predicate and a set of properties specific to that predicate.
+Each logical form induces natural language sentences that preserve its semantics.
 Logical Form Example Templates Example Sentences
@@
 entity=e)
-partwhole (whole_agent=$\hskip 1.00006pt\land\hskip 1.00006pt_{i=1}^{n}\texttt{a}_{i}$, [$\texttt{a}_{1}$], …, and [$\texttt{a}_{n}$] combine the [f]s that they have.
-Alice and Bob combine the fruits that they have.
-part_agent1=a1,
-part_agentn=an,
-whole_entity=f,
-part_entity=e)
+partwhole (whole_agent=$\hskip 1.00006pt\land\hskip 1.00006pt_{i=1}^{n}\texttt{a}_{i}$, [$\texttt{a}_{1}$], …, and [$\texttt{a}_{n}$] have [q] [e]s combined.
+quantity=q, Alice and Bob have 8 fruits combined.
+whole_entity=e)
 comp-eq (agentA=a, The number of [e]s that [c] has more than [d] is equal to the difference between the number of [e]s that [a] and [b] have.
@@
 entity=e)
-3 A Formal Treatment of Math Word Problems
-We aim to generate new problems of arbitrary complexity, so we require a formalism for characterizing complexity in a precise manner.
-First, we explain how the semantics expressed in math word problems can be written as sequences of logical forms (§ 3.1).
-We then show how to apply inference rules on these logical forms to deduce new ones (§ 3.2).
-This leads to a view of problem solutions as proof trees, the structure of which can be used to characterize the reasoning required to solve the problem.
 3.1 Math word problems as logical forms
-Each math word problem is represented as a sequence of logical forms under the formalism from Opedal et al.
-(2023; 2024).
-A logical form is a statement about a truth in the world that is being described by the problem, representing some arithmetic relationship between the number of entities possessed by one or several agents.
+Each MWP is represented as a sequence of logical forms under the formalism from Opedal et al.
+(2023; 2024), which is a subset of first-order logic.
+A logical form is a truth statement about the world that is being described by the problem, representing some arithmetic relationship between the number of entities possessed by one or several agents.
 Such statements can refer to the number of entities an agent has, how many more entities an agent has relative to another, the action of an agent giving their entities to another agent, etc.
-Formally, a logical form represents the semantics of a single sentence, and consists of a predicate that takes a set of properties as arguments.
-Every logical form has an agent property, an entity property and a quantity property.
-Other, optional, properties include unit and attribute.
-We define a world model as a sequence of logical forms that align with the sentences of a problem text, thus, representing the semantics of the entire problem.
-See Fig.
-5 for the world model representation of the problem from Fig.
-1.
+Formally, a logical form represents the semantics of a single sentence and consists of a predicate that takes a set of properties as arguments.
+Every logical form has an agent property, an entity property, and a quantity property.
+Other, optional properties include unit and attribute.
+We define a world model as a sequence of logical forms that align with the sentences of a problem text, thus representing the semantics of the whole problem.
 For a given predicate and set of properties, we write predicate(property1=x, property2=y, ...) to express a logical form.
 Property names are, however, often omitted for brevity, as in predicate(x, y, ...).
-There are two types of predicates: (i) those that represent ownership over some entity by an agent, called container (abbreviated cont), and (ii) those that represent arithmetic relationships between two sets of properties.
+There are two types of predicates: (i) those that represent the ownership of some entity by an agent, called container (abbreviated cont), and (ii) those that represent arithmetic relationships between two sets of properties.
 The relationship predicates correspond to arithmetic concepts and are based on a taxonomy from the learning sciences (Riley et al., 1983); we use transfer , comparison (abbreviated comp), partwhole and comp-eq.
 [^1] The transfer and part-whole concepts are usually called change and combine, respectively, in learning science literature (Nesher et al., 1982).
-The comp-eq predicate was introduced in this paper in order to construct nonlinear problems of arbitrary length.
-Table 1 gives concrete example sentences for logical forms with each of these predicates.
-In words, a cont logical form expresses the fact that an agent owns a particular entity in some quantity (e.g., “Alice has 5 apples”).
-A comp logical form expresses how much more some agent owns of a particular entity than some other agent (e.g., “Bob has 3 fewer apples than Alice”).
-A transfer logical form expresses the transfer of a certain quantity of an entity from one agent to another (e.g., “Alice gave Bob 3 apples”).
-A partwhole logical form expresses a superset relation between the combined entities of several agents (e.g., Alice and Bob combine their fruits).
-Finally, a comp-eq logical form expresses that the quantity of a particular comp between two agents is equal to that of a comp between two other agents (e.g., “The difference between the number of apples that Alice and Bob have is the same as the difference between the number of apples that Charlie and David have”).
-The logical forms in the world model of a math problem are separated into the body and question.
-The question is a single logical form which appears at the end of the world model.
-It is characterized by having a placeholder variable instead of an explicit quantity, which corresponds to the correct answer of the problem.
+The comp-eq predicate was introduced in this paper in order to construct nonlinear problems of arbitrary length—see App.
+D for details.
+Table 1 gives concrete example sentences for logical forms with each of these predicates, Fig.
+5 shows the world model representation of the problem in Fig.
+1, and App.
+A provides further intuition.
+[^2] In order to denote that the items of several agents are combined under a partwhole, we treat the conjunction of agents $\hskip 1.00006pt\land\hskip 1.00006pt_{i=1}^{n}\texttt{a}_{i}=\texttt{a}$ as another agent in Table 1.
+The same is done for entities, as shown in § 3.1.
+The logical forms in the world model of a problem are separated into body and question.
+The question is a single logical form that appears at the end of the world model.
+It holds a placeholder variable instead of an explicit quantity, which represents the correct answer to the problem.
 For example, the interrogative sentence “How many apples does Alice have?” is represented by the question cont(Alice, q, apples), where q is a placeholder variable.
-The body consists of all logical forms in the world model which are not questions.
-The answer to a problem is the logical form declared by the question with the correct numerical quantity substituted into its placeholder variable, although we will sometimes refer to only the correct numerical quantity as the answer as well.
+The body consists of all logical forms in the world model that are not questions.
+The answer to a problem is the logical form declared by the question with the correct numerical quantity substituted into its placeholder variable.
 Table 2: Inference rule templates (left) with corresponding examples of rule applications in natural language (right).
@@
 Attribute and unit properties are excluded here, but analogous rules exist in which they are present.
-All inference rules are commutative except the transfer rule; see footnote 3.
+All inference rules are commutative except the transfer rule; see footnote 4.
 Rule applications of an inference rule with premises $L_{1},L_{2},\dots,L_{N}$ and consequent $L$ are written as $L^{\prime}_{1}.L^{\prime}_{2}.\,\dots\,.L^{\prime}_{N}.\vdash L^{\prime}.$, where $L^{\prime}$ is a natural language expression for the logical form $L$.
@@
 3.2 Chain-of-thought solutions as proof trees
-Solving math word problems is a form of deductive reasoning, where the correct answer follows necessarily from what is stated in the text through a combination of rules of arithmetic and world knowledge.
-We use such derivations, called proofs, to characterize the structure of problems.
-The proofs in our setup are analogous to those found in other deductive systems, e.g., of first-order logic.author=andreas,color=blue!40,size=,fancyline,caption=,]This could be a good place to cite other NLP papers which have formalized reasoning through deduction systems.
-Any suggestions?
-abu:there is a lot we could cite here; maybe going all the way back to Montague, surveys of formal semantics, categorial grammar (where parsing itself is formalized as a reasoning problem)
+Solving MWPs is a form of deductive reasoning, where the correct answer follows from what is stated in the text through a combination of rules of arithmetic and world knowledge.
+We use such derivations, called proofs, to characterize these problems.
+The proofs in our setup are analogous to those in other proof systems, such as natural deduction Gentzen (1935); Pfenning (2004).
+[^3] In fact, each of our inference rules shown in § 3.1 can be decomposed into the primitive inference rules of natural deduction, using first-order logic expressions.
+App.
+B demonstrates by example how this is done.
 […]
-A math word problem can be solved by applying inference rules until the answer of the problem has been proved.
-Formally, a proof tree is a rooted ordered tree where each node is labeled with a logical form $L$, which is the conclusion of a proof step with premises $L_{1},\dots,L_{N}$, matching the labels of the child nodes in the same order.
+A math word problem can be solved by applying inference rules until the answer to the problem has been proved.
+Formally, a proof tree is a rooted ordered tree where each node is labeled with a unique logical form $L$, which is the conclusion of a proof step with premises $L_{1},\dots,L_{N}$, matching the labels of the child nodes in the same order.
 Leaf nodes are thus labeled with axioms.
-A proof of a particular problem is a proof tree where the labels on the leaf nodes are logical forms in the body of the problem, and the label at the root is the answer to the problem.
+A proof of a particular problem is a proof tree where the labels on the leaf nodes are the logical forms in the body of the problem, and the label at the root is the answer to the problem.
 The tree shown in Fig.
 1 illustrates a proof of the problem given in the same figure.
-Proof complexity.
+Complexity of proof trees.
 We seek a systematic way to analyze whether LLMs can generalize from simple to more complex proofs.
-The definition of proof given above gives rise to several ways of characterizing reasoning, out of which we consider four: linearity, depth, width, and ordering.
-These are defined below.
-We say that a proof tree is linear if every one of its inference rules takes at most one premise that is not an axiom, and nonlinear otherwise.
-In other words, every step in the solution to a linear problem only uses at most one fact not directly present in the problem.
-The proof tree given in Fig.
-1 is nonlinear because the proof step that deduces the answer uses two premises that are not themselves axioms.author=andreas,color=blue!40,size=,fancyline,caption=,]We should give an example of a linear problem as well so that this is easier to understand abu: agreed, maybe in appendix?
-The depth of a proof tree is defined as its height, i.e., the maximum path length from the root node to any leaf node.
+[^5] Note that our use of the term “complexity” is different from its conventional use in proof theory, in which proof complexity refers to the efficiency of a proof system (Cook & Reckhow, 1979).
+The definition of proof given above gives rise to several ways of characterizing reasoning, out of which we consider four: depth, width, shape, and ordering.
+The depth of a proof tree is defined as its height, i.e.
+the maximum path length from the root node to any leaf node.
 Next, the width of a proof tree is the number of leaf nodes, or axioms, that it contains.
@@
 1 thus has depth 2 and width 5.
-In words, this means that the answer is two proof steps “away” from any of the axioms, and that the problem has five (declarative) sentences.
+In words, this means that the answer is two proof steps “away” from any of the axioms and that the problem has five (declarative) sentences.
 [^6] In this paper, the width of the proof tree matches the number of declarative sentences (i.e., all sentences except the question) in the corresponding problem since we restrict the logical forms and sentences to be one-to-one.
-In general, that need not be the case since a single sentence could be written as a conjunction of several logical forms.
-One could also imagine problems that have irrelevant sentences, which do not represent any logical form at all.
-Lastly, we define the canonical ordering to be the ordering of leaf nodes as visited left-to-right.
+In general, this need not be the case since a single sentence could be written as a conjunction of several logical forms.
+One could also construct problems that have irrelevant sentences which either represent logical forms that will not be used in the proof or have no corresponding logical form at all.
+We distinguish two forms of proof-tree shapes: We say that a proof tree is linear if each of its proof steps takes at most one premise that is not an axiom, and nonlinear otherwise.
+[^7] We note that our notion of linear proofs is distinct from linear logic (Girard, 1987).
+In other words, every step in the solution to a linear problem uses only at most one fact not directly present in the problem; see Fig.
+6 for an example.
+The proof tree given in Fig.
+1 is nonlinear because the proof step that deduces the answer uses two premises that are not themselves axioms.
+Lastly, we define canonical ordering as the ordering of the leaf nodes when visited from left to right.
 Note that the sentences in the natural language annotation from Fig.
-1 follows the canonical ordering of the proof tree.
-Informally, this is the easy way of ordering the sentences in the problem, since it matches the ordering of the proof steps.author=andreas,color=blue!40,size=,fancyline,caption=,]Would be great if we could find support for this statement in cogsci literature However, there exist multiple alternative orderings in general.
-In particular, given a proof tree that only contains commutative inference rules (see footnote 3), all orderings of the leaf nodes are valid.
+1 follow the canonical ordering of the proof tree.
+Informally, this is an easy way to order the sentences in the problem, since it matches the ordering of the proof steps.
+However, there exist multiple alternative orderings in general.
+In particular, given a proof tree that only contains commutative inference rules (see footnote 4), all orderings of the leaf nodes are valid.
 For instance, swapping the ordering of the first two leaf nodes in the tree in Fig.
@@
 4 Evaluation Framework
-Equipped with the formalism explained in § 3, we propose an evaluation framework which tests LLMs on MWPs of arbitrary proof complexity, called MathGAP.
+Equipped with the formalism explained in § 3, we propose an evaluation framework that tests LLMs on MWPs that have arbitrarily complex proofs, called MathGAP.
 In § 4.1, we explain the generation method behind MathGAP.
@@
 We propose a method for synthetic generation of problems and their CoT solution explanations.
-It consists of three high-level steps: (i) sample a proof tree, (ii) map the logical forms at the leaf nodes into a natural language problem, and (iii) map the proof steps into a CoT solution.
-[^5] Jin et al.
-(2023) presents a related method for generating causal reasoning problems.
+The approach consists of three high-level steps: (i) sample a proof tree, (ii) map the logical forms at the leaf nodes into a natural language problem, and (iii) map the proof steps into a CoT solution.
+[^8] Our pipeline is similar to Opedal et al.’s (2024), except that we use a top-down procedure to generate full proof trees, rather than a sequential one to generate only axioms.
+This is a more general approach that enables generation of proofs with arbitrary shapes, e.g., nonlinear ones using comp-eq ; see App.
+D.
+Our pipeline is also similar to the method proposed by Jin et al.
+(2023) for generating causal reasoning problems.
 Fig.
 1 gives an illustration of these steps; note that the agent properties are abbreviated for brevity.
-In detail, the procedure is as follows: We first sample a logical form at the root of the proof tree.
-We then sample a specific inference rule as well as its premises.
-We recursively repeat this procedure for each premise until a predetermined stopping criterion has been reached, resulting in a proof tree.
+In detail, the procedure is as follows: We first sample a logical form to label the root of the proof tree.
+We then sample an inference rule whose conclusion matches the root, and use it to connect the root node to child nodes representing the premises of the sampled inference rule.
+We repeat this procedure recursively for each premise at the leaf nodes of the current tree until a predetermined stopping criterion has been reached, resulting in a proof tree.
 The stopping criterion is user-specified; it could be, e.g., when the tree has reached a certain depth or width.
-The categorical properties for the logical forms (i.e., agent, entity, attribute, and unit) are sampled uniformly at random from handwritten vocabularies.
-The quantities are sampled uniformly at random from a user-specified range; in our study presented in § 5 we use $2{-}20$.
-We then traverse the leaves of the proof tree in some order and convert each into a sentence via a natural language template, forming the text corresponding to the body of the problem.
+The categorical properties for the logical forms (i.e., agent, entity, attribute, and unit) are sampled uniformly at random from some vocabularies.
+We then convert each leaf node into a sentence via a natural language template in some order, forming the text corresponding to the body of the problem.
 The default order follows the proof tree’s canonical ordering, but other orders are possible as well.
 The sentences are sampled uniformly at random from a predefined set of templates; some example templates are given in Table 1.
-The question is converted from the logical form at the root of the proof tree, and is sampled from a set of interrogative sentences.
+The question is converted from the logical form at the root of the proof tree and is sampled from a set of interrogative sentences.
 In the final step, we generate the CoT annotation for the problem.
@@
 [^9] Note that the choice of traversing the nodes in post order corresponds to applying each inference rule as soon as all of its premises are available.
-In addition, post-order traversal visits the axioms in the canonical ordering.
-Alternative orders of tree traversal would violate at least one of these constraints.author=andreas,color=blue!40,size=,fancyline,caption=,]If we want to comment on similarity to the ordering of human reasoning etc we should look into literature on that.
-Start with the figural effect reference by Eisape et al.
-(2024) and the proof ordering references by Boguraev et al.
-(2024) We then convert each proof step into a sentence via a natural language template.
-The templates used for the axioms are identical to those used in the body of the problem text.
-In other words, the proof steps corresponding to axioms simply repeat the relevant part of the problem text.
-For the other proof steps, we have templates which are designed to explain the computation step that is performed.
+In addition, post-order traversal visits the axioms in canonical ordering.
+Alternative orders of tree traversal would violate at least one of these constraints.
+We then convert each proof step into a sentence via a natural language template.
+The templates used for the axioms are identical to those used in the body of the problem text; that is, the proof steps corresponding to axioms simply repeat the relevant part of the problem text.
+For the other proof steps, we use templates designed to explain the computation step that is performed.
 Fig.
-1 illustrates the mapping between nodes in the proof tree and sentences in the CoT annotation.
+1 illustrates the mapping between the nodes in the proof tree and the sentences in the CoT annotation.
 For instance, the third sentence in Fig.
 1 explains the conclusion cont(Lucy, 27, apple), which is deduced from the two leftmost nodes in the tree: “So Lucy has $17+10=27$ apples”.
-Our pipeline is similar to the one proposed by Opedal et al.
-(2024), except that we use a recursive procedure to generate each proof step starting at the root, instead of a sequential one.
-This is a more general approach which also enables generation of nonlinear proofs.
-[^7] The comp-eq rule enables the generation of nonlinear proof trees with arbitrary depth and is not included in the list of relations from Opedal et al.
-(2023).author=andreas,color=blue!40,size=,fancyline,caption=,]We need to explain the nuance of building nonlinear proofs in an appendix.
-We can create a nonlinear ones of depth two with only comp , but for deeper than that we need comp-eq.
-In the experiments we sample depth 2 examples of both comp and comp-eq.
 4.2 MathGAP: Mathematical Generalization on Arithmetic Proofs
-With the generation method just described, we can generate new synthetic problems where we control for the structure of the proof and its complexity.
-MathGAP can be used for various forms of systematic studies on arithmetic reasoning; here, we focus on OOD generalization.
-Specifically, we can select a family of problems of interest (e.g., linear problems under comp inference rules) and generate a train-test split of problems for which the problems in the test set are more complex than those in the training set (e.g., containing problems with deeper proofs).
-The generation method can be flexibly adapted to fit specific studies.
+With our generation method we can generate new synthetic problems where we control for the structure of the proof and its complexity.
+This mitigates the risk that the problems are contaminated.
+MathGAP can be used for various forms of systematic studies on arithmetic reasoning; here, we focus on in-context easy-to-hard OOD generalization.
+Specifically, we select a family of problems of interest (e.g., linear problems under comp inference rules) and generate a train-test split of problems in which problems in the test set are more complex than those in the training set (e.g., requiring deeper proofs).
+The generation method can be flexibly adapted.
 One can add new logical forms and inference rules, restrict the choice of existing ones during sampling, or create logical forms for irrelevant sentences (Shi et al., 2023).
-One can also vary the text templates for the logical forms and inference rules, the vocabularies for agents, entities, attributes and units, and the range of numbers from which the quantities are sampled.
-In addition, one may increase linguistic diversityauthor=andreas,color=blue!40,size=,fancyline,caption=,]There must be a better, more linguistically correct term for this by incorporating a last step in which a language model paraphrases the templated texts, or, alternatively, use a procedure like the one proposed by Mishra et al.
-(2024).
-In our study we do not perform such paraphrasing, since we want to guarantee that the sentences are faithful to the proof trees from which they are generated.
-However, future studies using MathGAP may choose to prioritize differently.
-5 Experiments
-author=andreas,color=blue!40,size=,fancyline,caption=,]Remind the reader somewhere here of why our evaluation framework is neat.
-We could have done an analysis on, say, GSM8k, but we prefer our synthetic data so that we can (i) avoid data contamination and (ii) control the proof tree characteristics
+One can also vary the text templates for the logical forms and inference rules, the vocabularies for agents, entities, attributes, and units, and the range of numbers from which the quantities are sampled.
+In addition, one may increase lexical and syntactic diversity by using an LLM to paraphrase the templated texts, e.g., through a procedure like Mishra et al.’s (2024).
+We do not perform such paraphrasing for the data in our experiments since it could introduce bias (Boguraev et al., 2024; Opedal et al., 2024) and produce problems that are unfaithful to their formal specifications.
+5 Generalizing to Complex Proofs
 We apply MathGAP to study how LLMs generalize from proof demonstrations of simple problems in context to solve problems of higher complexity at inference.
-We test OOD generalization through several sets of experiments: depth generalization for linear problems (§ 5.1), width generalization for linear problems (§ 5.2), depth generalization for nonlinear problems (§ 5.3), and generalization to permutations (§ 5.4).
-We are also interested in how the choice of the distribution of problems provided in context affects performance, which is often not obvious (Min et al., 2022; Turpin et al., 2023).
+We test OOD generalization through several sets of experiments: depth generalization for linear problems (§ 5.1), width generalization for linear problems (§ 5.2), depth generalization for nonlinear problems (§ 5.3), and generalization to permutations of the axioms (§ 5.4).
+Our public code repository provides further problem variations.
 Experimental setup.
-For each set of experiments, the general experimental setup is as follows: We generate five test sets of different degrees of complexity with 400 problems in each.
-We then generate model predictions for the problems in these test sets under four different prompts: (i) A zero-shot baseline prompt, which contains only the test problem; (ii) A prompt of primitive examples, in which all problems contain only one proof step of the same inference rule that is used in the test problem; (iii) A prompt of examples within a range of complexities, all of which are simpler than the test problem;author=abu,color=orange!40,size=,fancyline,caption=,]maybe we should format these settings differently so that it’s clearer that we will refer to them later in this section; either by bolding or maybe by naming them “Prompt (i):” (iv) An in-distribution baseline prompt, containing examples that are of the same complexity as the test problem.
-[…]
-Responses are generated using greedy decoding and a maximum length of 4,096 tokens.
-Model predictions are obtained by extracting the last number occurring in the model’s generated output.
+For each set of experiments, the general experimental setup is as follows: We generate multiple test sets of different degrees of complexity with 400 problems in each.
+We then generate model predictions for the problems in these test sets under four different prompts: (i) A zero-shot baseline prompt, which contains only the test problem; (ii) A prompt of primitive examples, in which all problems contain only one proof step of the same inference rule that is used in the test problem; (iii) A prompt of examples within a range of complexities, all of which are simpler than the test problem; (iv) An in-distribution baseline prompt, containing examples that are of the same complexity as the test problem.
+Note that prompts (i) through (iii) evaluate in-context OOD generalization, since the test examples are from a distribution that is different from that of the in-context examples.
+For prompts (ii) through (iv) we generate a new set of in-context examples for every test problem.
+This is done in order to avoid bias in the results towards any particular set of examples.
+We set the number of examples to 12, except for the experiments on nonlinear problems for which we use 5—see App.
+C for the rationale.
+Importantly, the in-context problems are provided together with their CoT solution annotations, as generated through the MathGAP pipeline.
+For prompt (iii), we make sure that there is at least one problem of every complexity in the range.
+This prompt can be viewed as a form of in-context curriculum learning (Bengio et al., 2009), [^10] Related approaches have been explored in previous work (Li et al., 2022; Liu et al., 2024b).
+except that we randomize the ordering of the examples rather than arranging them in increasing order of complexity.
+This is done since in-context orderings can have a large effect on performance (Lu et al., 2022), and we do not want our results to overfit to any particular ordering.
+More details on the data and prompt are given in App.
+C.
+Responses are generated using greedy decoding and a maximum context length of 4,096 tokens.
+Model predictions are obtained by extracting the last number occurring in the model output.
 We report answer accuracies and compute 95% confidence intervals using bootstrap sampling (Efron, 1992).
-We evaluate the following models on all test sets: Mixtral-8x7B (Jiang et al., 2024), Llama3 with 8B and 70B parameters (Llama Team, 2024), GPT-3.5 Turbo and GPT-4o (OpenAI, 2024).
-[^10] The maximum context windows for these models are 65k, 8k, 8k, 16k, 128k, respectively.
-Combined with a maximum number of output tokens of 4,096, the models have a theoretically sufficient context size to solve all test problems.
-The largest test problems (nonlinear depth 6) have around 2.5k tokens, including CoT annotation.
-For the OpenAI models we used “gpt-3.5-turbo-0125” and “gpt-4o-2024-05-13” as the model ids for all experiments.
-Figure 2: Answer accuracies for generalization to increasing depth and width for linear problems, across models and in-context distributions.
+We evaluate the following models on all test sets: Mixtral-8x7B (Jiang et al., 2024a), Llama3 with 8B and 70B parameters (Llama Team, 2024), GPT-3.5 Turbo and GPT-4o (OpenAI, 2024).
+Figure 2: Answer accuracies for generalization to increasing depth and width for linear problems across models and in-context distributions.
 Depth is increased using inference rules involving comp and transfer .
 Width is increased using partwhole.
+See App.
+E for example problems.
 5.1 Linear Depth Generalization
 Problem sets.
-We consider linear problems using both comp and transfer .author=andreas,color=blue!40,size=,fancyline,caption=,]We should give example proof trees for all our settings in appendix The five test sets consist of problems with depths between $6{-}10$.
+We consider linear problems using both comp and transfer .
+The five test sets consist of problems with depths between $6{-}10$.
 As such, the examples provided in the range of setting (iii) described above have depths between $1{-}5$.
-We give an example of a problem with depth $6$ in § B.1.
+We give an example of a problem with depth $6$ in § E.1.
 Results.
-We illustrate the results in the top and middle rows of Fig.
+The results are illustrated in the top and middle rows of Fig.
 2.
@@
 For the majority of cases, performance tends to be larger when providing a range of examples of varying complexity as compared to only primitive examples.
-The ranged OOD cases often show comparable performance to that of in-distribution context, suggesting that these models benefit from demonstrations on how problems of varying complexity are solved.
+The ranged OOD cases often show comparable performance to that of the in-distribution context, suggesting that these models benefit from demonstrations of problems with varying complexity.
 Llama3-8B is an outlier in that its solving accuracy for in-distribution context is (substantially) lower as compared to other prompts.
-When inspecting the model outputs, we see that the model often fails to answer the inference question, recognizing instead that it has been provided a sequence of in-context examples.
+When inspecting the outputs we see that the model often fails to answer the inference question, recognizing only that it has been provided a sequence of in-context examples.
 This could potentially be a consequence of how Llama3 is trained, using smaller context windows for the initial stages of pretraining (Llama Team, 2024).
-Finally, we note that GPT-4o and the larger Llama3 model show close to perfect performance across all test sets.
+This appears to be remedied by scale since Llama3 70B shows near perfect performance.
 5.2 Linear Width Generalization
 Problem sets.
-To test generalization to greater proof width, we use partwhole problems with a fixed depth of $1$.
-The five test sets have widths between $6{-}10$ and the width of the problems in setting (iii) described above have widths between $1{-}5$.
+We test generalization to larger proof width using partwhole problems with a fixed depth of $1$.
+The five test sets have widths between $7{-}11$ and the problems in setting (iii) described above have widths between $2{-}6$.
+Note that this matches the width for the problems in § 5.1; for those problems the width was equal to the depth plus one.
+§ E.2 gives an example problem.
 Results.
@@
 2.
+Apart from the GPT models, the performance is generally lower as compared to deep problems with the same number of sentences (§ 5.1), often with a sharper decreasing trend.
+Such is the case even for Llama3 70B, which performed almost perfectly in the linear depth setting.
+This suggests that it is more difficult to generalize in terms of width than in terms of depth, which is somewhat surprising considering that these problems only have one proof step.
 For GPT-3.5 and Mixtral-8x7B, we observe that providing in-context examples improves accuracy over zero-shot in most cases.
-Like for linear depth generalization, it seems that having a range of OOD examples of varying complexity is superior to only including primitive examples.
-Llama3-8B exhibits worse performance with in-distribution context here as well.
-In general for the Llama models, adding OOD in-context examples usually leads to lower performance over zero-shot.
-Apart from GPT-3.5-Turbo, the performance is generally lower as compared to a deep problem with the same number of sentences (§ 5.1), often with a sharper decreasing trend.
-Such is the case even for Llama3 70B, which performed almost perfectly in the linear depth setting.
-This would suggest that it is more difficult to generalize in terms of width than in terms of depth.
-Again, GPT-4o demonstrates near perfect performance across all test sets.
-Figure 3: Answer accuracies for generalization to increasing depth and width for nonlinear problems, across models and in-context distributions.
+As for linear depth generalization, it seems that having a range of OOD examples of varying complexity is superior to only including primitive examples.
+For the Llama models, adding OOD in-context examples often leads to lower performance over zero-shot.
+Figure 3: Answer accuracies for generalization to increasing depth and width for nonlinear problems across models and in-context distributions.
 Depth is increased using inference rules involving comp and comp-eq .
-The in-distribution contexts were too large to fit in the prompts for the Llama3 models and GPT-3.5 Turbo at some of the higher depths.
+The in-distribution contexts were too large to fit in context for the Llama3 models and GPT-3.5 Turbo at some of the higher depths.
 App.
-C shows additional results on o1-preview.
+F shows results on o1-preview and DeepSeek-R1.
 5.3 Nonlinear Depth Generalization
 Problem sets.
-We generate test datasets containing nonlinear problems with comp and comp-eq , having the same form as the problem in Fig.
+We generate test datasets containing nonlinear problems with comp and comp-eq which have a similar form as the problem in Fig.
 1.
-Another example is shown in § B.2.
+The procedure is explained in App.
+D and a more complex, depth $3$ example is given in § E.3.
 We generate a test set for each depth between $3{-}6$.
-Thus, the ranged OOD context setting contains examples of depths between $1{-}2$.author=andreas,color=blue!40,size=,fancyline,caption=,]Mention somewhere how depth interacts with width for these problem types
+Thus, the ranged OOD context setting contains examples of depths between $1{-}2$.
 Results.
@@
 Although we provided the same model id as in the other cases—which according to the documentation points to the original “gpt-4o-2024-05-13” snapshot—this particular test was performed at a later point in time.
-but for the deepest problems, their performance tends to zero as well.
+but for the deepest problems their performance tends to zero as well.
+Thus, the models are able to generalize to some extent but the error rate increases with complexity, suggesting that LLMs are noisy reasoners (Prabhakar et al., 2024).
 Moreover, it is worth noting that, for many cases, providing in-context examples does not significantly improve performance over zero-shot.
-On the other hand, for lower depths it is usually beneficial to provide in-distribution contexts.
-We perform an additional analysis on o1-preview in App.
-C.
+On the other hand, for lower depths, it is usually beneficial to provide in-distribution contexts.
+We perform an additional analysis on o1-preview and DeepSeek-R1 in App.
+F.
+Comparison to linear problems.
+Nonlinear problems are more difficult than the linear ones, even when controlling for the number of axioms given in the problem (i.e., its width).
+To see this, note that a nonlinear problem with depth 3 has 10 axioms (see App.
+D).
+This is the same number of axioms as the linear comparison and transfer problems (§ 5.1) with depth 9 and the linear part-whole problems (§ 5.2) with width 10.
+Comparing the results across those problem sets, we observe considerably lower accuracies for the nonlinear problems in most cases.
+We believe there are two main explanations for this: (i) Solving nonlinear problems requires keeping intermediate conclusions in memory while proving other intermediate conclusions, before being able to use them for further proof steps.
+It is thus less predictable where the relevant tokens will be located within the context window as compared to linear problems, in which intermediate conclusions are used immediately in a new proof step.
+(ii) Nonlinear problems use the comp-eq rule, which is most likely less frequent in the training set.
+Inspecting the model outputs we find support for both of these explanations—errors show up both when a previously deduced conclusion is needed as well as when a comp-eq is required in a proof step.
 5.4 Order Generalization
@@
 This has the effect of moving one of the sentences in the canonically-ordered problem text to the beginning of the problem.
-We characterize complexity as the distance of the movement, with the hypothesis that a greater movement from a canonical ordering constitutes a more difficult problem.author=andreas,color=blue!40,size=,fancyline,caption=,]Add a figure to the appendix where we show a linear proof tree and number the canonical orderings and give examples of movements we consider.
-Similar style as Fig.
-1 but less stuff in it.
+An illustration is shown in Fig.
+7 and § E.4 gives further examples.
+We characterize complexity as the distance of the movement, with the hypothesis that a greater movement from a canonical ordering constitutes a more difficult problem.
 We create five test sets with movement distances between $1{-}5$.
-We do not include a primitive OOD in-context distribution, the notion of a primitive problem is ill-defined in this case.
+We do not include a primitive OOD in-context distribution since the notion of a primitive problem is ill-defined in this case.
 The range prompt includes movements of all distances between $1{-}5$ except the one present in the test set.
-Figure 4: Answer accuracies for generalization to permutations, across models and in-context distributions.
-We measure complexity as the distance of the movement of a sentences moved to the beginning of the problem, as compared to its position in the canonical ordering of the problem.
+Figure 4: Answer accuracies for generalization to permutations across models and in-context distributions.
+We measure complexity as the distance of the movement of a sentence to the beginning of the problem, as compared to its position in the canonical ordering of the problem.
 Results.
@@
 First, note that all models show lower performance on perturbed problems as compared to non-perturbed ones (compare to Fig.
-2).
+2), suggesting that they are sensitive to simple perturbations of axiom orderings.
 Second, there is a nonlinear relationship between accuracy and distance of the movement, in contrast to the monotonically decreasing relation we expected.
@@
 In addition, including in-context examples gives a larger boost in performance for short and long movements, as compared to medium-length ones.
+Inspecting the model outputs, we observe that the models often make logical mistakes when the sentence that has been moved is needed for the subsequent proof steps.
+These kinds of mistakes appear to be more common than arithmetic errors.
 6 Conclusion
-In this paper, we introduced MathGAP, a framework for evaluating LLMs on math word problems of arbitrary arithmetic proof complexity.
+This paper introduced MathGAP, a framework for evaluating LLMs on math word problems with proofs that are arbitrarily complex.
 MathGAP can flexibly generate synthetic arithmetic word problems with controllable proof tree characteristics.
-Importantly, this enables studies of OOD generalization that avoid issues of data contamination, since a train-test split where the distributions are different can be created programmatically.
+Importantly, this enables studies of easy-to-hard OOD generalization that avoid issues of data contamination, since a train-test split where the distributions are different can be created programmatically.
 In particular, the test set can be made arbitrarily more complex than the training set.
-We apply MathGAP to study whether LLMs can learn from simple examples in context to solve more complex problems in a test set.
-All LLMs tested show a decrease in performance as proof complexity increases through depth and/or width.
-We also find that LLMs are sensitive to the order in which sentences are presented in a particular manner: A problem is harder if a sentence is moved from the middle of the problem, as compared to the beginning or end.
-We find no clear relationship between the distribution of in-context examples and performance; zero-shot or OOD contexts sometimes yield performance higher than or comparable to in-distribution contexts.
-Finally, we demonstrate that we can use MathGAP to construct problems that are challenging even for state-of-the-art reasoning models like OpenAI o1—see App.
-C.
-Limitations
-Although we make sure to sample from a wide range of agent names, entities, attributes and units, and include multiple templates for the predicates, future work may investigate how our findings generalize to distributions of problems with higher linguistic diversity.
-It is also beyond the scope of the present study to consider problem texts in languages other than English; however, it is straightforward to extend MathGAP to other languages.
-In addition, we cannot guarantee that the distributions of our generated problems are different from those of any potential internal data-generating mechanisms from Meta or OpenAI.
-However, we deem it to be unlikely that they use a similar formalism as we do here, and it gets increasingly unlikely that problems are similar as complexity increases.
-While the more complex nonlinear problems are indeed challenging even for the most capable models, we note that the main purpose of our study was not to create a challenging evaluation set for state-of-the-art LLMs.
-If that is the goal, one could create harder problems by combining several inference rules (e.g., including the rate concept; Opedal et al., 2023), designing a wider variety of proof tree topologies, and presenting the axioms in arbitrary orderings.
-It is also possible to create more syntactically complex templates and use numbers on which it has been shown to be more challenging for models to perform arithmetic (Razeghi et al., 2022).
-Finally, it is hard to make a direct comparison between the difficulty of linear and nonlinear problems, since they use different inference rules.
-In particular, nonlinear proofs contain comp-eq predicates, which are not present in linear proofs.
+We applied MathGAP to study whether LLMs can learn from simple examples in context to solve more complex problems in a test set.
+All models showed a decrease in performance as complexity of proof trees increases through depth and/or width.
+The LLMs struggled more with wider problems than with deeper problems, and were worse at solving problems with nonlinear shapes as compared to linear ones.
+We also found that LLMs are sensitive to the order in which sentences are presented.
+We further demonstrated that we can use MathGAP to construct problems that are challenging even for state-of-the-art reasoning models like OpenAI o1 and DeepSeek-R1—see App.
+F.
+Code to generate problems with MathGAP, including problem types beyond those considered in this paper, can be found in our public code repository.
+Finally, see App.
+G for a discussion on limitations of the present work.
 Acknowledgments
-We thank Alessandro Stolfo, Shehzaad Dhuliawala, Yanick Zengaffinen, and Yilmazcan Ozyurt for useful feedback and discussion.
+We thank Alessandro Stolfo, Ryan Cotterell, Shehzaad Dhuliawala, Yanick Zengaffinen, and Yilmazcan Ozyurt for useful discussion.
 Andreas Opedal acknowledges funding from the Max Planck ETH Center for Learning Systems.
-Appendix A Example of a World Model
-Fig. 5 gives an example problem text along with its world model.
+Appendix A Intuition on Logical Forms
 Figure 5: World model (right) of a math word problem (left).
@@
 The sentences and logical forms labeled (1) through (5) are the body of the problem, (6) is the question and (7) is the answer.
-Appendix B Example Problems
-author=yilmazcan,color=white!40,size=,fancyline,caption=,]For each subsection, maybe a small comment about why this problem is depth 6 or 3, or width 6 etc?
+A cont logical form expresses the fact that an agent owns a particular entity in some quantity (e.g., “Alice has 5 apples”).
+A comp logical form expresses how much more some agent owns of a particular entity than some other agent (e.g., “Bob has 3 fewer apples than Alice”).
+A transfer logical form expresses the transfer of a certain quantity of an entity from one agent to another (e.g., “Alice gave Bob 3 apples”).
+A partwhole logical form expresses a superset relation between the combined entities of several agents (e.g., Alice and Bob combine their fruits).
+Finally, a comp-eq logical form expresses that the quantity of a particular comp between two agents is equal to the difference of the quantities of the entity of two other agents (e.g., “The number of apples that Alice has more than Bob is the same as the difference between the number of apples that Charlie and David have”).
+Fig. 5 gives an example problem text along with its world model.
+Appendix B Defining Inference Rules with Natural Deduction
+[…]
+  1. 1.
+    $\exists v_{1}(\texttt{Owner}(v_{1},\texttt{a})\land\texttt{Measure}(v_{1},\texttt{q}_{1})\land\forall x\in v_{1}.\texttt{e}(x))$ by Axiom (1st premise, definition of container as in Opedal et al.
+2023, App.
+D.2.1).
+  2. 2.
+    $\texttt{Owner}(v_{1},\texttt{a})\land\texttt{Measure}(v_{1},\texttt{q}_{1})\land\forall x\in v_{1}.\texttt{e}(x)$ by $\exists$-elimination [1].
+  3. 3.
+    $\exists e_{1}(\texttt{ComparisonAdd}(e_{1})\land\texttt{Source}(e_{1},v_{1})\land\texttt{Target}(e_{1},v_{2})\land\texttt{Time}(v_{1})=\texttt{Time}(v_{2})\land\exists r(\texttt{Arg}(e_{1},r)\land\texttt{Measure}(r,\texttt{q}_{2})\land\forall x\in r.\texttt{e}(x)))$ by Axiom (2nd premise, definition of comparison as in Opedal et al.
+2023, App.
+D.2.2).
+  4. 4.
+    $\texttt{ComparisonAdd}(e_{1})\land\texttt{Source}(e_{1},v_{1})\land\texttt{Target}(e_{1},v_{2})\land\texttt{Time}(v_{1})=\texttt{Time}(v_{2})\land\exists r(\texttt{Arg}(e_{1},r)\land\texttt{Measure}(r,\texttt{q}_{2})\land\forall x\in r.\texttt{e}(x))$ by $\exists$-elimination [3].
+  5. 5.
+    $\exists r(\texttt{Arg}(e_{1},r)\land\texttt{Measure}(r,\texttt{q}_{2})\land\forall x\in r.\texttt{e}(x))$ by $\land$-elimination [4].
+  6. 6.
+    $\texttt{Arg}(e_{1},r)\land\texttt{Measure}(r,\texttt{q}_{2})\land\forall x\in r.\texttt{e}(x)$ by $\exists$-elimination [5].
+  7. 7.
+    $\texttt{Arg}(e_{1},r)\land\forall x\in r.\texttt{e}(x)$ by $\land$-elimination [6].
+  8. 8.
+    $\texttt{ComparisonAdd}(e_{1})\land\texttt{Arg}(e_{1},r)\land\forall x\in r.\texttt{e}(x)$ by $\land$-introduction [7].
+  9. 9.
+    $\forall e\forall r(\texttt{ComparisonAdd}(e)\land\texttt{Arg}(e,r)\land\forall x\in r.\texttt{e}(x)\to\exists v\exists o\exists q(\texttt{Target}(e,v)\land\texttt{Owner}(v,o)\land\texttt{Measure}(v,q)\land\forall x\in v.\texttt{e}(x))$ by Axiom (theorem on the existence of containers from existing relations as in Opedal et al.
+2023, App.
+D.4).
+  10. 10.
+    $\texttt{ComparisonAdd}(e_{1})\land\texttt{Arg}(e_{1},r)\land\forall x\in r.\texttt{e}(x)\to\exists v\exists o\exists q(\texttt{Target}(e_{1},v)\land\texttt{Owner}(v,o)\land\texttt{Measure}(v,q)\land\forall x\in v.\texttt{e}(x))$ by $\forall$-elimination [9].
+  11. 11.
+    $\exists v\exists o\exists q(\texttt{Target}(e_{1},v)\land\texttt{Owner}(v,o)\land\texttt{Measure}(v,q)\land\forall x\in v.\texttt{e}(x))$ by $\to$-elimination [8, 10].
+  12. 12.
+    $\texttt{Target}(e_{1},v_{2})\land\texttt{Owner}(v_{2},\texttt{b})\land\texttt{Measure}(v_{2},\texttt{q}_{t})\land\forall x\in v_{2}.\texttt{e}(x)$ by $\exists$-elimination [11].
+  13. 13.
+    $\texttt{Owner}(v_{2},\texttt{b})\land\texttt{Measure}(v_{2},\texttt{q}_{t})\land\forall x\in v_{2}.\texttt{e}(x)$ by $\land$-elimination [12].
+  14. 14.
+    $\forall e\forall v_{s}\forall v_{t}\forall m_{s}\forall m_{t}\forall r(\texttt{ComparisonAdd}(e)\land\texttt{Arg}(e,r)\land\texttt{Source}(e,v_{s})\land\texttt{Target}(e,v_{t})\land\texttt{Measure}(v_{s},m_{s})\land\texttt{Measure}(v_{t},m_{t})\to m_{s}+r=m_{t})$ by Axiom (theorem defining the semantics of comparison as in Opedal et al.
+2023, App.
+D.3).
+  15. 15.
+    $\texttt{ComparisonAdd}(e_{1})\land\texttt{Arg}(e_{1},r)\land\texttt{Source}(e_{1},v_{1})\land\texttt{Target}(e_{1},v_{2})\land\texttt{Measure}(v_{1},\texttt{q}_{1})\land\texttt{Measure}(v_{2},\texttt{q}_{t})\to\texttt{q}_{1}+\texttt{q}_{2}=\texttt{q}_{t}$ by $\forall$-elimination [14].
+  16. 16.
+    $\texttt{Measure}(v_{1},\texttt{q}_{1})$ by $\land$-elimination [2].
+  17. 17.
+    $\texttt{Measure}(v_{2},\texttt{q}_{t})$ by $\land$-elimination [13].
+  18. 18.
+    $\texttt{ComparisonAdd}(e_{1})\land\texttt{Arg}(e_{1},r)$ by $\land$-elimination [8].
+  19. 19.
+    $\texttt{Source}(e_{1},v_{1})\land\texttt{Target}(e_{1},v_{2})$ by $\land$-elimination [4].
+  20. 20.
+    $\texttt{ComparisonAdd}(e_{1})\land\texttt{Arg}(e_{1},r)\land\texttt{Source}(e_{1},v_{1})\land\texttt{Target}(e_{1},v_{2})\land\texttt{Measure}(v_{1},\texttt{q}_{1})\land\texttt{Measure}(v_{2},\texttt{q}_{t})$ by $\land$-introduction [4, 16, 17, 18, 19].
+  21. 21.
+    $\texttt{q}_{1}+\texttt{q}_{2}=\texttt{q}_{t}$ by $\to$-elimination [15, 20].
+  22. 22.
+    $\texttt{Owner}(v_{2},\texttt{b})\land\texttt{Measure}(v_{2},\texttt{q}_{1}+\texttt{q}_{2})\land\forall x\in v_{2}.\texttt{e}(x)$ by $=$-elimination [13, 21].
+  23. 23.
+    $\exists v_{2}(\texttt{Owner}(v_{2},\texttt{b})\land\texttt{Measure}(v_{2},\texttt{q}_{1}+\texttt{q}_{2})\land\forall x\in v_{2}.\texttt{e}(x))$ by $\exists$-introduction [22].
+Note that the conclusion (step 22) is the first-order logic equivalent of cont(b, q1 + q2, e), which is the conclusion of the comparison inference rule.
+Appendix C More Details on Experiments
+Generated problems.
+The quantities are sampled uniformly at random from the range $2{-}20$.
+We use a handwritten list of 52 English-language names for most problems, with the exception of the larger nonlinear problems for which we use a larger list of 4,945 predominantly English-language names from https://github.com/dominictarr/random-name.
+For the entities we use a handwritten word list with 51 items.
+We restrict each problem to have one entity, We determine uniformly at random whether the problem has attributes, units, or neither of them.
+Prompts.
+[…]
+Models.
+The maximum context windows for Mixtral-8x7B, Llama3 8B, Llama3 70B, GPT-3.5 Turbo and GPT-4o are 65k, 8k, 8k, 16k, 128k, respectively.
+Combined with a maximum number of output tokens of 4,096, the models have a theoretically sufficient context size to solve all test problems.
+The largest test problems (nonlinear depth 6) have around 2.5k tokens, including CoT annotation.
+For GPT-3.5 Turbo and GPT-4o we used “gpt-3.5-turbo-0125” and “gpt-4o-2024-05-13” as the model ids for all experiments.
+Appendix D Nonlinear Comparison Problems
+In this section we explain why we have introduced the comp-eq predicate to generate nonlinear trees. We also give the procedure for generating the nonlinear problems used for the experiments in § 5.3.
+Motivating example.
+[…]
+Now, say we are following the top-down generation procedure described in § 4.1 and wish to expand the tree to higher depths.
+Note that the two leaf nodes labeled with comp are impossible to expand further.
+For instance, the comp(B, A, 3, a) would be deduced by the two premises cont(A, 5, a) and cont(B, 8, a), but those two logical forms are already present elsewhere in the tree and are thus not available.
+This leaves only the two cont nodes for further expansion.
+However, any further subtrees of the cont nodes will be linear, since none of the comp nodes generated along the way will be able to be expanded.
+[^13] This reasoning assumes that we only use the inference rules with cont and comp from § 3.1.
+Deeper nonlinear problems could be achieved by using the partwhole rule as well.
+[…]
+Procedure.
+We want to sample a nonlinear proof tree with depth $D$.
+We start by sampling either a cont node or a comp node uniformly at random to determine the root of the tree.
+Next, we sample an inference rule for which the root logical form is the conclusion.
+We then sample inference rules from the child nodes in breadth-first order until all leaves except the ones labeled with comp-eq (which cannot be expanded) have depth $D$.
+For comp nodes, there is only one inference rule possible (the third row in § 3.1).
+For cont nodes we sample a comp rule (i.e., the first row in § 3.1) if the cont node has depth $D-1$, and a comp-eq rule (i.e., the fifth row in § 3.1) otherwise.
+This ensures that we can continue to grow the tree nonlinearly until the specified depth, following the reasoning given in the paragraph above.
+Note that the width of the proof trees generated with this procedure will be the number of leaf nodes in a perfect binary tree with depth $D$, i.e., $2^{D}$, plus the number of nodes labeled with comp-eq .
+Thus, a depth 2 problem will have a width of either 4 or 5, a depth 3 problem will have a width of 10, and a depth 4 problem will have a width of either 20 or 21.
+In general, the width $W(D)$ for a proof tree of depth $D$ is:
+$W(D)=2^{D}+\sum_{d=0}^{D-2}f_{c}(d),$
+where $f_{c}(d)$ gives the number of comp-eq nodes as a function of the level $d$, following the recursion
+$f_{c}(0)=\begin{cases}1\quad\text{if root is {cont} missing}\\ 0\quad\text{if root is {comp} missing}\end{cases}$
+$f_{c}(1)=\begin{cases}1\quad\text{if root is {cont} missing}\\ 2\quad\text{if root is {comp} missing}\end{cases}$
+$f_{c}(d)=f_{c}(d-1)+2f_{c}(d-2),d\geq 2.$
+Appendix E Example Problems and Illustrations
+Figure 6: Example of a linear problem using comp with depth 3 and width 4.
+This problem is linear because each parent node in the tree has at most one child that is not a leaf node.
+Compare to the nonlinear problem shown in Fig.
+1.
 In this section we provide example problems for each set of experiments (§§ 5.1, 5.2, 5.3 and 5.4).
-B.1 Example linear comparison problem (depth 6)
+E.1 Example linear comparison problem (depth 6)
 Jackson has 16 red bottles of soap.
@@
 What is the number of red bottles of soap that Mia has?
-B.2 Example nonlinear comparison problem (depth 3)
+E.2 Example partwhole width problem (width 7)
+Emily has 5 apples.
+Lily has 8 bananas.
+Abigail has 9 bananas.
+Benjamin has 11 grapes.
+Christopher has 20 apples.
+Mila has 16 grapes.
+Sophia has 11 watermelons.
+If everyone sums up the fruits that they have, how many fruits does everybody have in total?
+E.3 Example nonlinear comparison problem (depth 3)
 Ella has 11 yellow plates.
@@
 What is the number of yellow plates that Henry has?
-B.3 Example partwhole width problem (width 7)
-Emily has 5 apples.
-Lily has 8 bananas.
-Abigail has 9 bananas.
-Benjamin has 11 grapes.
-Christopher has 20 apples.
-Mila has 16 grapes.
-Sophia has 11 watermelons.
-If everyone sums up the fruits that they have, how many fruits does everybody have in total?
-B.4 Permuted problems
+E.4 Permuted problems
+Figure 7: The same proof tree as Fig.
+6 but with a permuted ordering as in the experiments discussed in § 5.4.
+In this specific example, the leaf node labeled with comp(L, J, 11, apple) is visited first (the other nodes follow post-order traversal).
+This translates into moving the sentence “Lucy has 11 more apples than John” to the front of the problem.
+As compared to the canonically ordered problem shown in Fig.
+6, the sentence has moved two steps.
 Canonically ordered problem text
@@
 [Original Position] How many computers does Logan have in their collection?
-Appendix C Additional Analysis on OpenAI o1
+Appendix F Additional Analysis on Reasoning Models
 Fig.
-6 shows additional results on the OpenAI o1-preview model, “o1-preview-2024-09-12”, evaluated on the nonlinear test sets.
-We only used zero-shot prompting due to the high costs of performing inference with o1-preview.
+8 shows additional results on the OpenAI o1-preview model (“o1-preview-2024-09-12”) and DeepSeek-R1 (DeepSeek-AI, 2025), evaluated on the nonlinear test sets.
+We only used zero-shot prompting due to the high costs of performing inference on these models.
 Using a simple, straightforward prompt is also consistent with OpenAI’s current recommendations.
-We see that the o1 model shows superior performance to all models from the main text.
+We see that these models show superior performance to all models from the main text, with o1 generalizing better than R1.
 However, as with the other models, the performance decreases as complexity increases.
-Importantly, we note that performance can be improved by increasing the limit on the number of output tokens.
+Importantly, we noted that performance can be drastically improved by increasing the limit on the number of output tokens.
 The results presented in Fig.
 8 used a token limit of 4,096, like the experiments in the main paper.
-We compared the effect of the token limit on an additional test set of 400 nonlinear problems with depth 7.
+For o1, we compared the effect of the token limit on an additional test set of 400 nonlinear problems with depth 7.
 With a token limit of 4,096, the model could only answer $0.25\%$ of the problems correctly.
 However, with a token limit of 10,000, it answered $76.5\%$ of the problems correctly.
-These results are noteworthy, and suggests that the new inference technique used by the o1 model is highly effective in decomposing complex problems into smaller steps.
-To test the model’s limits, we performed a final evaluation on a subset of 100 of the depth 7 problems where the sentences in each problem were randomly ordered.
-Recall from § 3.2 that permuting the sentences in random order is valid since all the inference rules used in these problems are commutative.
-In this experiment, we allowed 25,000 output tokens, which is the current recommendation by OpenAI.
-On this set, the answer accuracy was only $5\%$.
-Thus, we conclude that MathGAP is future-proof in the sense that it can generate problems on which current state-of-the-art models fail.
+For R1, the performance on problems with depth 6 is increased to $77.5\%$ when using a token limit of 10,000.
+These results are noteworthy and suggests that the inference techniques used by these models is highly effective for decomposing complex problems into smaller steps.
+Figure 8: Answer accuracies for generalization to increasing depth and width for nonlinear problems for OpenAI o1-preview and DeepSeek-R1 under zero-shot prompting and a maximum context length of 4,096 tokens.
+To test the limits of these models, we performed a final evaluation on a subset of 100 of the depth 7 problems where the sentences in each problem were randomly ordered (apart from the question).
+Recall from § 3.2 that permuting the sentences in random order is valid since all the inference rules used in our nonlinear problems are commutative.
+In this experiment, we allowed 25,000 output tokens for both models, which is the current recommendation by OpenAI.
+On this set, the answer accuracy was only $5.0\%$ for o1 and $11.0\%$ for R1.
+Thus, we conclude that MathGAP is future-proof in the sense that it can generate problems on which current state-of-the-art reasoning models fail.
 We stress, however, that the aim of this paper is not to construct a challenging test set for state-of-the-art models; there are several other ways that MathGAP can be used to create problems that might be even more difficult.
-Figure 6: Answer accuracies for generalization to increasing depth and width for nonlinear problems for OpenAI o1-preview under zero-shot prompting.
+Appendix G Limitations
+We did not prioritize linguistic diversity in our problem sets; future work may investigate how our findings generalize to distributions of problems with higher linguistic diversity.
+We did also not consider or study the possibility of token biases (Jiang et al., 2024b).
+It is beyond the scope of the present study to consider problem texts in languages other than English; however, it is straightforward to extend MathGAP to other languages.
+In addition, we cannot guarantee that the distributions of our generated problems are different from those of any potential internal data-generating mechanisms from Meta or OpenAI.
+However, we deem it to be unlikely that they use a similar formalism as we do here, and it gets increasingly unlikely that problems are similar as complexity increases.
+While the more complex nonlinear problems are indeed challenging even for the most capable models, we note the purpose of our study was not to create a challenging evaluation set for state-of-the-art LLMs.
+If that is the goal, one could create harder problems by combining several inference rules (e.g., additionally including the rate concept; Opedal et al., 2023), allowing a wider variety of proof-tree shapes, and presenting the axioms in arbitrary orderings.
+It is also possible to create more syntactically complex templates (and paraphrase those using LLMs) and use numbers on which it is more challenging for models to perform arithmetic (Razeghi et al., 2022).
+Finally, it is hard to make a direct comparison between the difficulty of linear and nonlinear problems, since they use different inference rules.
+In particular, nonlinear proofs contain comp-eq predicates, which are not present in linear proofs.